### v1

# UFID: A Unified Framework for Input-level Backdoor Detection on Diffusion Models

Abstract
Diffusion Models are vulnerable to backdoor attacks, where malicious attackers inject backdoors by poisoning some parts of the training samples during the training stage. This poses a serious threat to the downstream users, who query the diffusion models through the API or directly download them from the internet. To mitigate the threat of backdoor attacks, there have been a plethora of investigations on backdoor detections. However, none of them designed a specialized backdoor detection method for diffusion models, rendering the area much under-explored. Moreover, these prior methods mainly focus on the traditional neural networks in the classification task, which cannot be adapted to the backdoor detections on the generative task easily. Additionally, most of the prior methods require white-box access to model weights and architectures, or the probability logits as additional information, which are not always practical. In this paper, we propose a Unified Framework for Input-level backdoor Detection (UFID) on the diffusion models, which is motivated by observations in the diffusion models and further validated with a theoretical causality analysis. Extensive experiments across different datasets on both conditional and unconditional diffusion models show that our method achieves a superb performance on detection effectiveness and run-time efficiency. The code is available at https://github.com/GuanZihan/official_UFID.
1 Introduction
Diffusion models [22, 40, 41, 38] have emerged as the new state-of-the-art family of generative models due to their superior performance [16] and wide applications across a variety of domains, ranging from computer vision [5, 8], natural language processing [2, 23, 28], and robust machine learning [7, 9]. Despite their success, the training of diffusion models consumes a tremendous amount of time and computational resources. Consequently, it is a common practice to directly utilize third-party models via an API or directly download them from the internet. The setting is referred to as Model-as-a-Service (MaaS).
Recently, it was found that diffusion models are vulnerable to backdoor attacks [11, 14, 15, 42, 20], where the malicious attackers poison some parts of the training samples with a predefined trigger pattern in the training stage. Consequently, the behavior of the diffusion models can be adversarially manipulated whenever the trigger pattern appears in the input query, while it remains normal with clean input queries. This vulnerability not only poses a serious threat to downstream users, such as inadvertently generating inappropriate images for children who query the model, but also poses risks to companies or artists with extensive copyright interests. For instance, it could manipulate the model to generate images that bypass copyright restrictions [43].
Figure 1: Visualization of the input-level backdoor detection setting.
To mitigate the threat of backdoor attacks, numerous investigations have been conducted focusing on backdoor detections [45, 31, 46]. However, none of these studies specifically aim to develop a detection method for generative models, leaving this area significantly under-explored. Moreover, most of these methods are confined to traditional neural networks in classification tasks and are not readily adaptable to generative tasks due to two primary challenges: ❶ More Diverse Failures: Unlike merely generating a fixed target image (e.g., a Hello-Kitty image), backdoored diffusion models can be manipulated to produce a specific class of images (e.g., cat images), or even images with a specified abstract concept (e.g., erotic images). This implies that the target images are not necessarily unique but can vary as long as they belong to the designated target class. Consequently, the poisoning process differs significantly from simply inducing a poisoned model to predict a fixed target label in image classification tasks. This variability substantially complicates detection in generative tasks. ❷ Multi-Modality Attack Surface: Diverging from traditional image classifiers, which involve a single modality, diffusion models (e.g., Stable Diffusion) are capable of supporting multiple modalities. This diversity necessitates a unified framework for backdoor detection in diffusion models.
Additionally, most of the prior methods require white-box access to the model weights and architecture, or the probability logits as additional information, which are not always practical in the MaaS setting. For the backdoor detection problem in the MaaS setting (i.e., black-box input-level backdoor detection [21, 30]; visualized in Figure 1), we assume no prior information about model weights and architectures, but that only the user query and the generated results from diffusion models are available. Intuitively, we aim to work as a firewall that receives users’ queries and determines whether the generated result can be returned to the users. In particular, we approve and forward the DNN’s prediction results for clean images while rejecting those for poisoned images. The challenges of the setting stem from two requirements: (1) efficiency: our method should not significantly influence the model’s inference speed, and (2) effectiveness: our method should effectively distinguish backdoor and clean samples. Therefore, how to detect backdoor samples in the MaaS setting for the diffusion models is far from a trivial question.
To address the above challenges, we propose a Unified Framework for Input-level backdoor Detection (UFID) on diffusion models. Our method is motivated and underpinned by a causality analysis where the strong backdoor attack serves as the confounder, introducing a spurious path from input to target images. Furthermore, the spurious path embedded in the diffusion model remains consistent when we perturb the input samples with Gaussian noise. We further validate the causal analysis with a theoretical analysis. Driven by the analysis, we develop our final method for backdoor detection in diffusion models. Specifically, for conditional diffusion models, we gradually perturb the inputs with different random noises; for unconditional diffusion models, we perturb the text inputs by augmenting them with additional public textual prompts. Based on our previous analysis, a backdoored input tends to produce similar images regardless of perturbations, while a clean sample generates multiple diversified images. This contrast in image generation allows us to use the similarity of the generated images as a basis for detecting backdoor samples. Extensive experiments across different datasets on both conditional and unconditional diffusion models demonstrate that our method achieves superb performance on detection effectiveness and run-time efficiency.
To sum up, our contributions in the work include: (1) First Unified Backdoor Detection Framework for Diffusion Models. To the best of our knowledge, our work is the first unified framework for detecting backdoor samples in diffusion models; (2) Novel Causality Analysis and Theoretical Analysis. We are the first paper to apply causality analysis for analyzing backdoor attacks on generative tasks; Besides, we also provide a theoretical analysis to validate the effectiveness of our method; (3) SOTA performance. Extensive results on various datasets empirically show that our method achieves outstanding performance in terms of detection effectiveness and runtime efficiency.
2 Preliminaries
2.1 Diffusion Models
Without loss of generality, diffusion models are composed of two components: (1) Diffusion Process: a data distribution $q(x)$ is diffused to a target distribution $r(x)$ within T timestamps. (2) Training Process: a diffusion model with parameter $\theta$ is trained to align with the reversed diffusion process, i.e., $p_{\theta}(x_{i-1}|x_{i})=\mathcal{N}(x_{i-1};\mu_{\theta}(x_{i}),\sigma_{\theta}(x_{i}))=q(x_{i-1}|x_{i})$. In this section, we introduce the basic version DDPM [22].
DDPM.
DDPM assumes the target distribution $r(x)=\mathcal{N}(0,I)$ and the diffusion process $q(x_{i}|x_{t-1})=\mathcal{N}(x_{t};\sqrt{1-\beta_{i}}x_{i-1},\beta_{i}I)$, where the $\{\beta_{i}\}_{i=1}^{T}$ is a pre-defined variance schedule that controls the step sizes. Furthermore, let $\alpha_{i}=1-\beta_{i}$ and $\bar{\alpha_{i}}=\Pi_{t=1}^{i}\alpha_{t}$. By minimizing the loss function $\|\epsilon-\epsilon_{\theta}(\sqrt{\bar{\alpha_{i}}}x_{0}+\sqrt{1-\bar{\alpha_{i}}}\epsilon,i)\|^{2}$, the diffusion model is expected to be able to correctly predict the added noise given the input $x_{i}$ at time $i$. In the inference stage, DDPM generates images by sampling from the Gaussian distribution $\mathcal{N}(0,I)$ from time $i=T$ to $i=0$ with the generative process $p_{\theta}(x_{i-1}|x_{i})=\mathcal{N}(x_{i-1};\mu_{\theta}(x_{i}),\sigma_{\theta}(x_{i}))$, where $\mu_{\theta}(x_{i})=\frac{1}{\alpha_{i}}(x_{i}-\frac{1-\alpha_{i}}{1-\bar{\alpha_{i}}\epsilon_{\theta}(x_{i},i)})$ and $\sigma_{\theta}(x_{i})=\frac{(1-\bar{\alpha}_{i-1})\beta_{i}}{1-\bar{\alpha}_{i}}$.
2.2 Backdoor Attacks on Diffusion Models
Different from launching backdoor attacks to the traditional models (e.g., classifiers [19, 12]), which could be achieved by poisoning training dataset, injecting backdoors into the diffusion models is much more complicated. A typical backdoor attack pipeline [11, 14, 15] on diffusion models consists of three steps: (1) the attackers first need to mathematically define the forward backdoor diffusion process, i.e., $x_{0}^{b}\rightarrow x^{b}_{T}$, where the $x_{0}^{b}$ denotes the target image and the $x_{T}^{b}$ denotes the trigger image; (2) then the attackers train the diffusion models to align with the backdoored reversed process; (3) in the inference stage, the diffusion models can be prompted to generate the target images when the input contains the trigger pattern, but behave normally when the input is clean (e.g., pure gaussian noise for the DDPM model).
2.3 Threat Model
We adopt a similar setting as in  [21, 30], which is named "black-box input-level backdoor detection". Specifically, two parties are considered in the setting: Attacker and Defender. We present a detailed description of each party as follows.
Attacker.
We assume a strong attacker that is capable of training a backdoored model and uploading the backdoored model to the public models market, such as cloud platforms, GitHub, and Hugging Face. The downstream users, such as software engineers, can then be authorized to deploy the model in the local end device. In the inference stage, the malicious attacker can activate the backdoor by querying the deployed model with an additional trigger in the input.
Defender.
The defender aims to conduct efficient and effective input-level black-box backdoor detection on the user side. Specifically, efficiency requires that the detection process does not significantly impact the response time of user queries, while effectiveness requires the detection process to distinguish backdoor samples and clean samples with a high accuracy rate. The defender is assumed to only have access to the deployed model’s outputs (e.g., the generated image from the diffusion model), without any prior information such as model weight and embedding.
Problem 1 (Input-level Backdoor Detection).
For a given diffusion model represented as $f_{\theta}(\cdot)$, the defender’s objective of the defender is to develop a detector, denoted by $\mathcal{A}(\cdot)$. This detector aims to output $\mathcal{A}(\bm{x}_{backdoor})=1$ for poisoned inputs, and $\mathcal{A}(\bm{x}_{clean})=0$ for clean inputs.
The challenge of Problem 1 arises from three factors:
  1. More Diverse Failures. Backdoored diffusion models can be triggered to generate specific classes of images (e.g., cat images), or even images with a specified abstract concept (e.g., erotic images), extending beyond fixed target labels in traditional classification tasks.
  2. Multi-Modality Attack Surface. Unlike traditional image classifiers that involve only one modality, diffusion models (e.g., Stable Diffusion) can support a variety of modalities.
  3. Limited Information & Efficiency Guarantee. The detector $\mathcal{A}(\cdot)$ has access only to the query images and the prediction labels returned by the DNN, and is expected to minimally impact the inference efficiency.
3 Methodology
Figure 2: Pipeline of our unified framework for backdoor detection on diffusion models.
3.1 Backdoor Attacks from a Causal View
To develop a backdoor detection algorithm for generative models, we first need to address a fundamental question: What distinguishes clean generation from backdoored generation, and how this distinction can be utilized in designing an effective detection algorithm? To this end, we propose to leverage causal inference to reveal the distinct mechanisms underlying clean and backdoored generation processes. Specifically, we present a causal graph that illustrates the comparison of the two processes in Figure 3. A causal graph is a directed acyclic graph that illustrates the causal relationships among variables, where each node represents a variable. An edge from a node $A$ to a node $B$ implies that the variable $A$ is the cause of the variable $B$ (denoted as $A\rightarrow B$). For simplicity, this figure only includes the unconditional diffusion model. However, this model can be easily extended to the conditional diffusion model by substituting the input image ($X_{T}$) with the input text ($T$).
Figure 3: Causal graph of clean and backdoored generation.
Clean Generation.
As depicted in Figure 3(a), the generated image ($x_{0}^{c}$) is dependent on the input noise image $x_{T}^{c}\sim\mathcal{N}(0,I)$. This relationship is termed the causal path, denoted as $X_{T}^{c}\rightarrow X_{0}^{c}$. Consequently, adding Gaussian noise $\epsilon\sim\mathcal{N}(0,I)$ to $x_{T}^{c}$ results in a new input ${x^{\prime}}_{T}^{c}=x_{T}^{c}+\epsilon=\mathcal{N}({x^{\prime}}_{T}^{c};0,2I)$, leading to a different generated image ${x^{\prime}}_{0}^{c}$, as shown in Figure 3(c).
Backdoored Generation.
As shown in Figure 3(b), a backdoor attack $A$ modifies an image $x_{T}$ by injecting a trigger $\delta$ and changing the image generation process towards the target image $x_{0}^{b}$, denoted as $X_{T}^{b}\leftarrow A\rightarrow X_{0}^{b}$, where $x_{T}^{b}=\delta+x_{T}$. This introduces a spurious path from $X_{T}^{b}$ to $X_{0}^{b}$, which lies outside the direct causal path $X_{T}^{b}\rightarrow X_{0}^{b}$, thereby establishing and strengthening erroneous correlations between the modified input noise and the target image. Consequently, generations from poisoned noise images are primarily influenced by this spurious path [17, 47, 29], while the direct causal path $X_{T}^{b}\rightarrow X_{0}^{b}$ plays a minor role, represented by a gray dotted line in Figure 3 (b). When an additional Gaussian noise $\epsilon\sim\mathcal{N}(0,I)$ is added to the backdoored noise image $x_{T}^{b}$, the new backdoored input becomes a Gaussian noise ${x^{\prime}}_{T}^{b}\sim\mathcal{N}(\delta,2I)$, which is a combination of a new noise image $x_{T}+\epsilon$ and the trigger pattern $\delta$. Since the trigger pattern remains in the new backdoored input, the spurious path continues to dominate the generation process. As a result, the generated image will remain unchanged.
In summary, for clean generation, a small perturbation significantly alters the output. However, triggers in backdoor samples tend to be robust features learned by neural network models. Consequently, minor perturbations of backdoor samples do not lead to substantial changes in the diffusion model’s generation results. The following theorems validate our insights from causal analysis.
Lemma 3.1.
[…]
The Lemma 3.1 states that two points sampled from a distribution with larger variance tend to be more diversified than those sampled from a distribution with lower variance. A detailed proof is provided in Appendix H.
Lemma 3.2.
Let $f_{\theta}$ and $f_{\tilde{\theta}}$ be two well-trained diffusion models as defined in the Assumption G.4. Let $x^{\prime}_{T}$ follow $\mathcal{N}(0,\rho^{2}I)$. Let the clean data distribution be $q(x)\sim\mathcal{N}(x_{c},\sigma_{c}I)$. We then have:
$\hat{x_{0}}=f_{\theta}(x^{\prime}_{T})\sim\mathcal{N}(x_{c},\frac{\sigma_{c}}{\rho^{2}}I)$ (1)
This Lemma 3.2 indicates that after adding Gaussian noise to the input, the mean of generated image keeps unchanged, while the variance is scaled by $\frac{1}{\rho^{2}}$. In our paper, $\rho^{2}=2$. A detailed proof is provided in Appendix I.
Theorem 3.3.
Given a diffusion model $f_{\theta}$, clean input noise $x_{T}\sim\mathcal{N}(0,I)$, and backdoor input noise $x_{T}^{b}=x_{T}+\delta\sim\mathcal{N}(\delta,I)$, if we perturb the clean input noise and backdoor input noise with some $\epsilon\sim\mathcal{N}(0,I)$ simultaneously, then the generated images from clean input noise will be much more diversified than that from backdoor input noise.
Proof.
The proof is under Lemma 3.1 and  3.2. We provide detailed proof in Appendix G.
3.2 Scenario 1: Unconditional Diffusion Models
[…]
$y_{i}^{j}={f_{\theta}(x_{i}^{j})},\forall x_{i}^{j}\in\mathbb{I},$ (2)
We further let $y_{i}=\{y_{i}^{j}|1\leq j\leq|\mathbb{M}|+1\}$ denote the generated batch. We can then determine whether the input query $x_{i}$ is a backdoored sample or not by inspecting $y_{i}$.
3.3 Scenario 2: Conditional Diffusion Models
Conditional diffusion models accept input from other modalities to guide the generation of user-intended images. For instance, in stable diffusion [36], textual input is used to query the model. However, it is evident that the detection approach employed for unconditional diffusion models cannot be directly applied to conditional diffusion models due to the inability to introduce Gaussian noise to discrete textual input. To adapt our detection method for conditional diffusion models, we have extended the detection approach with slight modifications, as depicted in the second row of Figure 2. We leverage variations in output diversity to distinguish between clean and backdoor samples. To further accentuate this diversity gap and enhance distinguishability in conditional setting, we propose appending the input text $x_{i}$ with a random phrase $ph_{j}$ selected from a diverse phrase pool containing completely distinct phrases, such as "Iron Man" and "Kitchen Dish Washer," denoted as =$\mathbb{N}=\{ph_{1},ph_{2},...ph_{|\mathbb{N}|}\}$, where $|\mathbb{M}|\ll|\mathbb{N}|$. For each input $x_{i}$, we repeat this process $|\mathbb{M}|$ times, resulting in augmented an input batch $\mathbb{I}=\{x_{i}\}\cup\{x_{i}^{j}|x_{i}^{j}=x_{i}\oplus ph_{j},\forall 1\leq j\leq|\mathbb{M}|\}$, where $\oplus$ denotes the string appending operator. Similarly, we can generate an image batch $y_{i}$ by querying the target stable diffusion model using Equation 2.
3.4 A Unified Framework for Backdoor Detection
As previously discussed, the diversity of images generated by a specific input $i$ can be leveraged to detect backdoors. To quantify this diversity, we employ a two-step process that involves calculating both pairwise and overall similarities within the generated batch. The detailed steps are as follows:
Pairwise Similarity Calculation.
Initially, we calculate semantic embeddings of generated images through a pre-trained image encoder (e.g., ViT-ImageNet [44] and CLIP [35]), denoted as $f_{\mathcal{E}}(\cdot)$. Subsequently, we calculate the local similarity for each pair of images in the generated batch using cosine similarity, represented by $S_{c}(\cdot,\cdot)$.
Graph Density Calculation.
Following this, we construct a weighted graph and compute its graph density to represent the overall similarity of all images within graph. Specifically, let $G_{i}=(V_{i},\mathcal{E}_{i})$ represent the similarity graph for input sample $x_{i}$, where $|V_{i}|=|\mathbb{M}|$ constitutes the set of vertices (symbolizing the generated images) and $\mathcal{E}_{i}$ is the set of edges. Each edge’s weight, connecting a pair of images $u,v\in V$, indicates their similarity score, denoted as $E[u,v]=S_{c}(u,v)$, where $E_{i}\in\mathbb{R}^{|\mathcal{E}|}$. In this similarity graph, the similarity between two generated images is interpreted as the distance within the graph. Next, we introduce graph density [4], as a novel metric for evaluating the overall similarity of the generated batch:
Definition 3.4.
The graph density $DS(G_{i})$ of the weighted similarity graph is defined as:
$DS(G_{i})=\frac{\sum_{(m<n)}S_{c}(f_{\mathcal{E}}(y_{i}^{m}),f_{\mathcal{E}}(y_{i}^{n}))}{|\mathbb{M}|(|\mathbb{M}|-1)}$
If $DS(G_{i})$ is greater than the threshold $\tau$, then it is determined as a backdoor sample, otherwise, it is a clean sample and the originally generated image shall be returned to the users. The total pipeline of our method is visualized in Figure 2 and the final detection algorithm is presented in Algorithm 1.
4 Experimental Results
4.1 Experimental Settings
Attack Baselines.
To the best of our knowledge, the existing backdoor attacks on diffusion models include two unconditional-DM-based backdoor attacks: TrojDiff [11] and BadDiffusion [14], and two conditional-DM-based backdoor attacks: Rickrolling [42] and Villandiffusion [15]. We consider all four backdoor attacks as our attack baselines. It is also noted that TrojDiff supports three types of attack objectives. Detailed descriptions of the four attack methods are provided in Appendix B.
Defense Baselines.
To the best of our knowledge, there are yet methods for detecting backdoor samples on diffusion models. Moreover, existing methods are built especially for classification models and therefore not applicable to the generative models.
Models and Datasets.
Different backdoor attacks are built based on different backbone models and samplers. To facilitate evaluation, TrojDiff [11] and BadDiffusion [15] are evaluated on DDPM [22], while VillanDiffusion [15] and Rickrolling [42] are evaluated on Stable Diffusion [36]. For the training datasets, we choose Cifar10 [26] and CelebA [32] for TrojDiff and BadDiffusion, and choose CelebA-HQ-Dialog [25] and Pokemon [34] for VillanDiffusion and Rickrolling.
Metrics.
Following the prior works on backdoor detection, we adopt three popular metrics for evaluating the effectiveness of our detection method: Precision (P), Recall (R), and Area under the Receiver Operating Characteristic (AUC).
Implementation Details
All the models are well-trained with the default hyper-parameters reported in the original papers so that they show a good performance in generating both clean images and backdoor images. Following the previous works [27, 21], we then evaluate our detection method with a positive (i.e., attacked) and a negative (i.e., clean) dataset. For evaluations against unconditional diffusion models, we randomly generate 1000 Gaussian noises as the clean queries (negative) and construct backdoor samples (positive) accordingly by blending the trigger pattern with the Gaussian noises. For evaluations against conditional diffusion models, we split the whole dataset into 90% train and 10% test following [15]. Then, we use the textual caption in the test subset as the clean queries (negative) and construct backdoor queries (positive) accordingly. The threshold value $\tau$ is determined by a small clean held-out validation dataset, where detailed descriptions are provided in Appendix C. The pre-trained encoder is set as CLIP-ViT-B32 [35], the magnitude size is set as 4, the poisoning rate is set as 10%, and the number of validation datasets is set as 20, by default. All the hyperparameters are evaluated in the ablation studies.
4.2 Main Results
Effectiveness
Table 1: Performance of the proposed detection method against backdoor attacks on unconditional diffusion models.
Dataset Backdoor Attacks Precision Recall AUC
Cifar10 TrojanDiff(D2I) 0.95 0.94 1.00
TrojanDiff(In-D2D) 0.93 0.93 0.98
TrojanDiff(Out-D2D) 0.93 0.92 1.00
BadDiffusion 0.93 0.95 1.00
Average 0.93 0.94 1.00
CelebaA TrojanDiff(D2I) 0.93 0.92 1.00
TrojanDiff(In-D2D) 0.90 0.89 0.96
TrojanDiff(Out-D2D) 0.91 0.92 0.98
BadDiffusion 0.97 0.95 1.00
Average 0.93 0.92 0.99
Table 2: Performance of the proposed detection method against backdoor attacks on conditional diffusion models.
Dataset Backdoor Attacks Precision Recall AUC
CelebaA-HQ-Dialog VillanDiffusion 0.92 0.95 0.96
Rickrolling 0.87 0.84 0.90
Average 0.90 0.90 0.93
Pokemon Caption VillanDiffusion 0.91 0.93 0.94
Rickrolling 0.83 0.85 0.91
Average 0.87 0.89 0.93
Table 2 presents the performance of our detection method against backdoor attacks on unconditional diffusion models, and Table 2 presents the performance on conditional diffusion models. As shown, the AUC values for different backdoor attack methods on all the evaluated datasets are over 0.9, suggesting that our method can effectively distinguish backdoor and clean samples.
Efficiency
Figure 4: Average inference speed of the proposed detection method against TrojDiff(D2I) on the Cifar10 dataset.
Figure 4 illustrates the efficiency of our detection method against TrojDiff(D2I) on the Cifar10 dataset. We query the diffusion models with 320 samples with a batch size of 64 and record the average inference speed, defined as the average time consumption for a sample. The y-axis comprises three components: ’Vanilla,’ representing the average inference speed without UFID; ’+augmented Batch Query’ representing the average inference speed after an augmented batch query; and ’+Similarity Calculation’ representing the average inference speed after similarity graph construction and calculation. Due to the increased number of query samples, UFID inevitably results in a lower inference speed than that of the vanilla mode. However, the increased time consumption is within expectations and will not significantly influence user experience in practice.
4.3 Ablation Studies
In this section, we discuss how the hyper-parameters influence the effectiveness of the detection method.
The Influence of Different Pre-trained Encoders.
Table 3: Performance of our detection method with different pre-trained encoders.
Encoder $\rightarrow$ ViT-ImageNet CLIP DINO V2
ViT-B ViT-L RN50 RN50x64 ViT-B ViT-L ViT-S ViT-B ViT-L
Precision 0.94 0.95 0.89 0.85 0.91 0.80 0.81 0.85 0.80
Recall 0.93 0.95 0.88 0.81 0.91 0.79 0.70 0.78 0.65
AUC 0.98 0.99 0.88 0.95 0.97 0.88 0.99 1.00 0.99
The pre-trained encoder is a crucial component in our detection method. To evaluate its impact on the effectiveness of our detection method, we test the performance of UFID when integrated with different pre-trained encoders in Table 3. For the space limit, we only present the results against TrojDiff(In-D2D) on the Cifar10 dataset in the main manuscript. More experiments are in the Appendix D. As shown, UFID works well when integrated with different pre-trained encoders. In particular, CLIP encoders show consistently good performance across different datasets, due to their strong generalization ability from the pre-training stage.
The Influence of Magnitude Set.
Figure 5: Performance with different sizes of magnitude set.
Figure 6: Performance with different amounts of available samples.
Figure 6 investigates the impact of the magnitude size on the running-time efficiency and effectiveness, where the X-axis denotes the magnitude size, the left y-axis denotes the performance values, and the right y-axis denotes the average inference speed. The first four backdoor attacks are evaluated on the Cifar10 dataset and the remaining two backdoor attacks are evaluated on the Pokemon dataset. As the figure shows, our detection method achieves a stably satisfactory performance against all backdoor attacks when the size is over four. Additionally, a size of four would also yield a balanced trade-off on efficiency and effectiveness.
The Influence of Available Validation Dataset.
Figure 6 presents the impact of the available validation dataset on the threshold values and the performance, where the X-axis values denote the number of available validation samples and the y-axis denotes the performance values. The first four backdoor attacks are evaluated on the Cifar10 dataset and the remaining two backdoor attacks are evaluated on the Pokemon dataset. As the figure suggests, with more validation samples available, the threshold values gradually increase, and the performances tend to become more stable. However, it is also noted that if the number of validation samples becomes exceedingly large, there is a slight drop in performance. A possible explanation for this phenomenon is that more validation samples also introduce more noisy information, leading to an unexpected threshold value.
The Influence of Different Poisoning Rate.
Figure 8 explores whether the performance of UFID is sensitive to the backdoor poisoning rate. We evaluate the UFID under poisoning rate from 0.05 to 0.30. The results reveal that our method can generally provide a satisfactory performance irrelevant to the poisoning rate. Moreover, with an increase in the poisoning rate, the performance becomes more stable.
Figure 7: Performance with different poisoning rates.
Figure 8: Evaluation of UFID against adaptive attacks.
Resilient against Adaptive Backdoor Attacks.
We evaluate our detection method against an adaptive attacker who already has prior information about our detection method. Therefore, the attacker might try to make the generated images more diversified to avoid being detected. Specifically, rather than training a diffusion model that maps the trigger to the target images (e.g., erotic images), the attacker maps the trigger to a target domain that contains both the target images and a small number of clean images. In this way, the attacker achieves a more stealthy backdoor attack by sacrificing the attack success rate. We further define the ratio between the number of clean images to the backdoor samples in this target domain as the "blending ratio". We evaluate the UFID against TrojDiff(D2I) and employ mean square error (MSE) between the generated backdoor images and the target image (e.g., Mickey Mouse) as the attack success rate. Figure 8 presents the performance of the UFID under different blending ratios. As shown, the performance of the UFID exhibits a slight decrease when the blending ratio rises. However, the average MSE across generated backdoor samples abruptly exceeds 0.15, which suggests the failure in injecting backdoors. The right-hand side also provides some samples in the Cifar10 dataset, where we notice that images with an MSE of 0.15 to the target image are already completely different from the target image.
4.4 Discussions
Visualizations of Similarity Graphs.
Figure 9: Similarity graphs against TrojDiff (In-D2D) attack on Cifar10 dataset.
To better understand how UFID helps to detect backdoor samples, we visualize the similarity graphs in Figure 9. Due to the space limit, we only present similarity graphs against the TrojDiff In-D2D attack on the Cifar10 dataset in the main manuscript. More qualitative examples are presented in the appendix E. Each node in the similarity graph denotes the generated images of the query batch, while each edge denotes the cosine similarity scores of the embedding of any two images. As shown, the similarity scores for the clean query batch are significantly lower than those for the backdoor query batch, validating our intuitions for backdoor detections.
Visualizations of Scores Distributions.
Figure 17 and Figure 18 present the distributions of the calculated scores $\mathcal{S}_{i}$ for backdoor samples and clean samples respectively. As shown, the distribution of backdoor samples tends to be more clustered in a narrow range, while that of clean samples tends to be spread out. Moreover, there is a distinct gap between the two distributions, suggesting that our method can effectively distinguish backdoor and clean samples.
5 Related Works
Diffusion Models.
Diffusion models have become a new state-of-the-art family of generative models in image synthesis by showing strong sample quality and diversity in image synthesis [22, 38]. Since then, diffusion models have been widely applied to different types of tasks, such as image generation [39, 41], image super-resolution [37, 6], and image editing [3, 13]. Moreover, researchers also found that the representations learned by diffusion models can also be used in other discriminative tasks such as segmentation [18] and anomaly detection [33].
Backdoor Attacks and Defenses on Diffusion Models.
Recently, inspired by the great success of diffusion models, a lot of works investigate the security vulnerabilities of diffusion models by launching backdoor attacks on diffusion models. From a high-level idea, malicious backdoor attackers aim to inject a special behavior in the diffusion process such that, once the predefined trigger pattern appears on the input, the special behaviors will be activated. To achieve this goal,  [11] proposed to add an additional backdoor injection task on the training stage and maliciously alter the sampling procedure with a correction term.  [14] proposed a novel attacking strategy by only modifying the training loss function.  [1] focused on launching attacks to text-to-image tasks, by injecting backdoors into the pre-trained text encoder.  [15] proposed a unified framework that covers all the popular schemes of diffusion models, including conditional diffusion models and unconditional diffusion models. However, backdoor defenses on diffusion models are highly under-explored. To the best of our knowledge, only  [1] investigated backdoor defenses on diffusion models. However, their work aims to detect whether a given model is backdoored, while our tasks focus on filtering backdoor samples for diffusion models in the inference stage, which are fundamentally different.
6 Conclusion and Future Directions
In this paper, we propose a simple unified framework for backdoor detection on diffusion models under the MaaS setting. Our framework is first motivated by a strict causality analysis on image generation and further validated by theoretical analysis. Motivated by the analysis, we design a unified method for distinguishing backdoor and clean samples for both conditional and unconditional diffusion models. Extensive experiments demonstrate the effectiveness of our method. Despite the great success, there are still many directions to be explored in the future. Firstly, our method relies heavily on a pre-trained encoder, which is not always practical. It is promising to design some model-free metrics to detect backdoor samples on diffusion models. Secondly, extending our unified framework to the newly emerging diffusion models is also of great interest.
Impact Statment
Diffusion Models have been widely adopted for generating high-quality images and videos. Therefore, inspecting the security of diffusion models is of great significance in practice. In this paper, we propose a simple unified framework that effectively detects backdoor samples for the diffusion models under a strict but practical scenario of Moel-as-a-Service (MaaS). As described in the threat model, our method is proposed from the perspective of a defender. Therefore, this paper has no ethical issues and will not introduce any additional security risks to diffusion models. However, it is noted that our method is only used for filtering backdoored testing samples but they do not reduce the intrinsic backdoor vulnerability of the deployed diffusion models. We will further improve our method in future works.
Appendix A Pseudo code of the UFID Detection Algorithm
The following pseudo code 1 presents the overall UFID detection algorithm.
Algorithm 1 Backdoor Detection on Diffusion Models
  Input: User Input $x_{i}$; Target Diffusion Model $f_{\theta}$; Detection threshold $\tau$.
  if unconditional model then
     $\mathbb{I}=\{x_{i}\}\cup\{x_{i}^{j}|x_{i}^{j}=x_{i}+\epsilon_{j},\forall 1\leq j\leq|\mathbb{M}|\}$
  else if conditional model then
     $\mathbb{I}=\{x_{i}\}\cup\{x_{i}^{j}|x_{i}^{j}=x_{i}+ph_{j},\forall 1\leq j\leq|\mathbb{M}|\}$
  end if
  for $j=1$ to $|\mathcal{M}|$ do
     $y_{i}^{j}={f_{\theta}(x_{i}^{j})},\forall x_{i}^{j}\in\mathbb{I}$
  end for
  $DS(G_{i})=\frac{\sum_{(m<n)}S_{c}(f_{\mathcal{E}}(y_{i}^{m}),f_{\mathcal{E}}(y_{i}^{n}))}{|\mathbb{M}|(|\mathbb{M}|-1)}$
  if $DS(G_{i})\leq\tau$ then
     Return the true generated image $y_{i}^{1}$.
  else
     Warning: $x^{i}$ is a backdoor query.
  end if
Appendix B More Details about Attack Baselines
TrojDiff [11].
We implement TrojDiff following the public code [^1] https://github.com/chenweixin107/TrojDiff on GitHub. As described, the TrojDiff framework encompasses three distinct types of backdoor attacks: D2I, In-D2D, and Out-D2D. D2I maps a pre-defined trigger to a specific target image; In-D2D associates the trigger with a specified class of images within the same distribution as the training datasets, and Out-D2D links the trigger to a specified class of images in a distribution different from the training dataset. Throughout all three backdoor attacks, a Hello Kitty image serves as the trigger pattern. Specifically, for D2I, we designate a Mickey Mouse image as the target image. In the case of In-D2D, the target class is chosen as the seventh class of the training dataset; for instance, on CIFAR-10, this corresponds to horse images (class 7). Similarly, in Out-D2D, the seventh class of the target dataset is selected as the target class. To illustrate, we opt for the MNIST dataset as our target dataset, where the seventh class consists of images of number seven. We give an illustration of TrojDiff in the Figure 10.
Figure 10: Examples of backdoor samples from TrojDiff.
BadDiffusion [14].
We implement BadDiffusion following the public code [^2] https://github.com/FrankCCCCC/baddiffusion_code/tree/master on GitHub. For BadDiffusion, we use an eye-glasses image as the trigger pattern, and the target image is a hat image used in the original paper. We give an illustration of BadDiffusion in the Figure 11.
Figure 11: Examples of backdoor samples from BadDiffusion.
VillanDiffusion [15].
We implement VillanDiffusion following the public code [^3] https://github.com/IBM/VillanDiffusion/tree/main on GitHub. As described, VillanDiffusion is a general framework for injecting backdoors into either conditional diffusion models or unconditional diffusion models. In this paper, we use VillanDiffusion specially refer to the backdoor attacks on conditional diffusion models. Specifically, the backdoor attacks is conducted over a pre-trained stable diffusion model [^4] https://huggingface.co/CompVis/stable-diffusion-v1-4/tree/main, so as to make the model generates target images once the caption trigger appears. We use "mignneko" as the caption trigger, and the Cat image used in the original paper as the target image, since these configurations are shown to perform well on different datasets in the original paper. We give an illustration of VillanDiffusion in the Figure 12.
Figure 12: Examples of backdoor samples from VillanDiffusion.
Rickrolling [42].
We implement Rickrolling following the public code [^5] https://github.com/LukasStruppek/Rickrolling-the-Artist on GitHub. As described, Rickrolling injects backdoors into the text encoder, by making the text encoder consistently generate the embedding of a target text when the trigger is present. It uses the Cyrillic o as the trigger and replaces $o$ in the original text to construct backdoor samples. The target text is chosen as "a drawing of a bird with blue eyes". We give a illustration of Rickrolling in the Figure 13.
Figure 13: Examples of backdoor samples from Rickrolling.
Appendix C More Details about How to Choose $\tau$.
Suppose we are given $n$ clean validation samples: $x_{1}$, $x_{2}$, …, $x_{n}$, then we take them as a batch and query the diffusion model as described in 2. In this way, a similarity graph $G$ can be constructed on this batch, with each edge denoting the similarity between any two generated images. Finally, for each node, we calculate an average similarity between this node to the other nodes. The maximal average value is used as the threshold $\tau$.
Appendix D More Ablation Studies on Pre-trained Encoders
In this section, we record performances of our detection method UFID against different backdoor attacks when integrated with different pre-trained encoders, where Table 4 is for TrojDiff(Out-D2D), Table 5 is for TrojDiff(D2I), Table 6 is for BadDiffusion, Table 7 is for VillanDiffusion, and Table 8 is for Rickrolling.
Table 4: Performance of our detection method against TrojDiff(Out-D2D) on the Cifar10 dataset with different pre-trained encoders.
Encoder $\rightarrow$ ViT-ImageNet CLIP DINO V2
ViT-B ViT-L RN50 RN50x64 ViT-B ViT-L ViT-S ViT-B ViT-L
Precision 0.93 0.94 0.95 0.85 0.90 0.83 0.90 0.88 0.80
Recall 0.92 0.93 0.94 0.78 0.88 0.75 0.87 0.84 0.67
AUC 1.00 1.00 1.00 0.99 1.00 1.00 1.00 1.00 1.00
Table 5: Performance of our detection method against TrojDiff(D2I) on the Cifar10 dataset with different pre-trained encoders.
Encoder $\rightarrow$ ViT-ImageNet CLIP DINO V2
ViT-B ViT-L RN50 RN50x64 ViT-B ViT-L ViT-S ViT-B ViT-L
Precision 0.94 0.93 0.95 0.83 0.90 0.84 0.90 0.88 0.80
Recall 0.93 0.92 0.95 0.74 0.87 0.76 0.88 0.85 0.68
AUC 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00
Table 6: Performance of our detection method against BadDiffusion on the Cifar10 dataset with different pre-trained encoders.
Encoder $\rightarrow$ ViT-ImageNet CLIP DINO V2
ViT-B ViT-L RN50 RN50x64 ViT-B ViT-L ViT-S ViT-B ViT-L
Precision 0.95 0.94 0.93 0.85 0.91 0.85 0.92 0.87 0.82
Recall 0.92 0.90 0.92 0.76 0.86 0.76 0.82 0.88 0.71
AUC 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00
Table 7: Performance of our detection method against VillanDiffusion on the Pokemon dataset with different pre-trained encoders.
Encoder $\rightarrow$ ViT-ImageNet CLIP DINO V2
ViT-B ViT-L RN50 RN50x64 ViT-B ViT-L ViT-S ViT-B ViT-L
Precision 0.62 0.63 0.94 0.89 0.91 0..88 0.84 0.83 0.85
Recall 0.57 0.67 0.95 0.92 0.93 0.89 0.88 0.82 0.88
AUC 0.64 0.68 0.97 0.94 0.94 0.90 0.91 0.90 0.92
Table 8: Performance of our detection method against Rickrolling on the Pokemon dataset with different pre-trained encoders.
Encoder $\rightarrow$ ViT-ImageNet CLIP DINO V2
ViT-B ViT-L RN50 RN50x64 ViT-B ViT-L ViT-S ViT-B ViT-L
Precision 0.51 0.62 0.90 0.81 0.83 0.77 0.75 0.80 0.76
Recall 0.55 0.64 0.89 0.81 0.85 0.76 0.75 0.78 0.75
AUC 0.66 0.68 0.94 0.90 0.91 0.89 0.86 0.87 0.84
Appendix E More Details about Similarity Graphs
We provide additional qualitative examples of similarity graphs in Figure 14, Figure 15 and Figure 16. Specifically, Figure 14 presents similarity graphs for backdoor attacks on the Cifar10 dataset, where the leftmost image represents a similarity graph for a clean query. Moving from left to right, we present qualitative examples of similarity graphs for backdoor queries under TrojDiff(D2I), TrojDiff(Out-D2D), and BadDiffusion, respectively. Moreover, Figure 15 and Figure 16 present similarity graphs for VillanDiffusion and Rickrolling backdoor attacks, where the left image represents a similarity graph for a clean query, and the right image is a similarity graph for a backdoor query.
Figure 14: Similarity graphs generated for backdoor attacks on the Cifar10 dataset. The leftmost image represents a similarity graph for a clean query. Moving from left to right, we present qualitative examples of similarity graphs for backdoor queries under TrojDiff(D2I), TrojDiff(Out-D2D), and BadDiffusion, respectively.
Figure 15: Similarity graphs generated for VillanDiffusion backdoor attacks. The left image represents a similarity graph for a clean query, and the right image is a similarity graph for a backdoor query.
Figure 16: Similarity graphs generated for Rickrolling backdoor attacks. The left image represents a similarity graph for a clean query, and the right image is a similarity graph for a backdoor query.
Appendix F More Details about Score Distributions
In Figure 17, we provide distributions of graph density scores $DS(G)$ for both clean and backdoor samples on Cifar10 dataset against TrojDiff(D2I), TrojDiff(Out-D2D), TrojDiff(In-D2D), and BadDiffusion attack, where the red bars denote the scores for clean samples, and the blue bars denote the scores for backdoor samples. Similarly, we provide distributions of graph density scores on the Pokemon dataset against VillanDiffusion and Rickrolling in Figure 18. For all of the distributions, we can notice there exists an obvious gap between the score distributions for backdoor samples and those for clean samples, suggesting that our detection method can effectively distinguish backdoor and clean samples.
Figure 17: Distributions of detection scores for backdoor samples and clean samples against unconditional diffusion models.
Figure 18: Distributions of detection scores for backdoor samples and clean samples against conditional diffusion models.
Appendix G Proof of Theorem 3.3
Theorem G.1.
Given diffusion model $f_{\theta}$, clean input noise $x_{T}\sim\mathcal{N}(0,I)$, and backdoor input noise $x_{T}^{b}=x_{T}+\delta\sim\mathcal{N}(\delta,I)$, if we perturb the clean input noise and backdoor input noise with some $\epsilon\sim\mathcal{N}(0,I)$ simultaneously, then the generated images from clean input noise will be more diversified than that from backdoor input noise.
Proof.
We begin our proof by first introducing the basic diffusion process for clean samples.
Definition G.2 (Clean Forward process).
Let $x_{0}\sim q(x)$ denote a sample from the clean data distribution, $x_{T}\sim\mathcal{N}(0,I)$ denote the pure Gaussian noise. Given the variance schedule $\{\beta_{t}\}_{t=1}^{T}$ in DDPM [22], define the forward process to diffuse $x_{0}$ to $x_{T}$ for clean samples:
$q(x_{t}|x_{t-1})=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}I)$ (3)
$q(x_{t}|x_{0})=\mathcal{N}(x_{t};\sqrt{\bar{\alpha}_{t}}x_{0},(1-\bar{\alpha}_{t})I),$ (4)
where $\alpha_{t}=1-\beta_{t}$ and $\bar{\alpha}_{t}=\Pi_{i=1}^{t}\alpha_{i}$.
After obtaining the forward process, then the diffusion model $f_{\theta}$ with parameter $\theta$ is trained to align with the reversed diffusion process, i.e., $p_{\theta}(x_{i-1}|x_{i})=\mathcal{N}(x_{i-1};\mu_{\theta}(x_{i}),\sigma_{\theta}(x_{i}))=q(x_{i-1}|x_{i})$, to learn how to obtain a clean image from a noise image. Here, we give the definition of the reverse process of clean samples:
Definition G.3 (Clean Reverse process).
The reverse process for clean samples is
$q(x_{t-1}|x_{t})=\mathcal{N}(x_{t-1};\mu_{\theta}(x_{t},t),\Sigma_{\theta}(x_{t},t)),$ (5)
$\mu_{\theta}(x_{t},t))=\frac{1}{\sqrt{\alpha_{t}}}(x_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}(x_{t},t)),$ (6)
$\Sigma_{\theta}(x_{t},t)=\frac{(1-\bar{\alpha}_{t-1})\beta_{t}}{1-\bar{\alpha}_{t}},$ (7)
Detailed proof can be found in [24]. In our setting, we assume that the attacker is able to deploy a well-trained diffusion model on the internet. Accordingly, we make the following assumptions:
Assumption G.4.
Assume a well-trained clean diffusion model $f_{\theta}$, designed to generate clean samples $x_{o}\sim q(x)$ from pure Gaussian noise $x_{T}\sim\mathcal{N}(0,I)$. Besides, we also assume there exists another well-trained diffusion model $f_{\tilde{\theta}}$ with parameters $\tilde{\theta}$, aimed at denoising $x^{\prime}_{T}=x_{T}+\epsilon=\mathcal{N}(x^{\prime}_{T};0,\rho^{2}I)$ back to the same clean data distribution $q(x)$ as that of $f_{\theta}$. The variance schedules $\{\beta_{t}\}_{t=1}^{T}$ for both models are identical.
This assumption implies that the noise predictors $\epsilon_{\theta}$ and $\epsilon_{\tilde{\theta}}$ are well-trained to accurately estimate the noise required to derive $x_{t}$ and $x^{\prime}_{t}$, respectively. As a result, both $f_{\theta}$ and $f_{\tilde{\theta}}$ can generate images following clean data distribution $q(x)$, given inputs following $\mathcal{N}(0,I)$ and $\mathcal{N}(0,\rho^{2}I)$, respectively. The forward and backward processes of $f_{\theta}$ are already defined from Equation 3 to 7. Notably, in our analysis, $\rho^{2}$ is set to 2 for clean samples to account for the addition of Gaussian noise. Hence, for $f_{\tilde{\theta}}$, the forward process is:
$q(x^{\prime}_{t}|x^{\prime}_{t-1})=\mathcal{N}(x^{\prime}_{t};\sqrt{1-\beta_{t}}x^{\prime}_{t-1},\beta_{t}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho^{2}}I)$ (8)
$q(x^{\prime}_{t}|x_{0})=\mathcal{N}(x^{\prime}_{t};\sqrt{\bar{\alpha}_{t}}x_{0},(1-\bar{\alpha}_{t}){\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho^{2}}I),$ (9)
With this diffusion process, $q(x)$ could be diffused to $\mathcal{N}(x_{T}^{\prime};0,\rho^{2}I)$ in $T$ steps. Then the $f_{\tilde{\theta}}$ aims to learn a generative process, such that $p_{\tilde{\theta}}(x^{\prime}_{t-1}|x^{\prime}_{t})=q(x^{\prime}_{t-1}|x^{\prime}_{t})$, which is,
$\begin{split}q(x^{\prime}_{t-1}|x^{\prime}_{t},x_{0})&=q(x^{\prime}_{t}|x^{\prime}_{t-1},x_{0})\frac{q(x^{\prime}_{t-1}|x_{0})}{q(x^{\prime}_{t}|x_{0})}\\ &\propto\exp\Big{(}-\frac{1}{2}\big{(}\frac{(x^{\prime}_{t}-\sqrt{\alpha_{t}}x^{\prime}_{t-1})^{2}}{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho^{2}}\beta_{t}}+\frac{(x^{\prime}_{t-1}-\sqrt{\bar{\alpha}_{t-1}}x_{0})^{2}}{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho^{2}}(1-\bar{\alpha}_{t-1})}-\frac{(x^{\prime}_{t}-\sqrt{\bar{\alpha}_{t}}x_{0})^{2}}{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho^{2}}(1-\bar{\alpha}_{t})}\big{)}\Big{)}\\ &=\exp\Big{(}-\frac{1}{2}\big{(}\frac{{x^{\prime}}_{t}^{2}-2\sqrt{\alpha_{t}}x_{t}x_{t-1}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{+\alpha_{t}}{x^{\prime}}_{t-1}^{2}}{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho^{2}}\beta_{t}}+\frac{{x^{\prime}}_{t-1}^{2}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{-2\sqrt{\bar{\alpha}_{t-1}}x_{0}}x^{\prime}_{t-1}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{+\bar{\alpha}_{t-1}x_{0}^{2}}}{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho^{2}}(1-\bar{\alpha}_{t-1})}-\frac{(x^{\prime}_{t}-\sqrt{\bar{\alpha}_{t}}x_{0})^{2}}{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho^{2}}(1-\bar{\alpha}_{t})}\big{)}\Big{)}\\ &=\exp\Big{(}-\frac{1}{2{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho^{2}}}\big{(}(\frac{\alpha_{t}}{\beta_{t}}+\frac{1}{1-\bar{\alpha}_{t-1}}){x^{\prime}}_{t-1}^{2}-(\frac{2\sqrt{\alpha_{t}}}{\beta_{t}}x^{\prime}_{t}+\frac{2\sqrt{\bar{\alpha}_{t-1}}}{1-\bar{\alpha}_{t-1}}x_{0})x^{\prime}_{t-1}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{+C(x^{\prime}_{t},x^{\prime}_{0})\big{)}\Big{)}}\\ &:=\mathcal{N}(x^{\prime}_{t-1};\mu_{\tilde{\theta}}(x^{\prime}_{t},t),\Sigma_{\tilde{\theta}}(x^{\prime}_{t},t)),\end{split}$ (10)
Following the standard Gaussian density function, the mean and variance can be parameterized as follows.
$\begin{split}\Sigma_{\tilde{\theta}}(x^{\prime}_{t},t)&=1/{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho^{2}}(\frac{\alpha_{t}}{\beta_{t}}+\frac{1}{1-\bar{\alpha}_{t-1}})=1/(\frac{\alpha_{t}-\bar{\alpha}_{t}+\beta_{t}}{\beta_{t}(1-\bar{\alpha}_{t-1})})=\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\cdot\beta_{t}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho^{2}}\\ \mu_{\tilde{\theta}}(x^{\prime}_{t},t)&=\frac{1}{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho^{2}}}(\frac{\sqrt{\alpha_{t}}}{\beta_{t}}x^{\prime}_{t}+\frac{\sqrt{\bar{\alpha}_{t-1}}}{1-\bar{\alpha}_{t-1}}x_{0})/\frac{1}{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho^{2}}}(\frac{\alpha_{t}}{\beta_{t}}+\frac{1}{1-\bar{\alpha}_{t-1}})\\ &=(\frac{\sqrt{\alpha_{t}}}{\beta_{t}}x^{\prime}_{t}+\frac{\sqrt{\bar{\alpha}_{t-1}}}{1-\bar{\alpha}_{t-1}}x_{0})\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\cdot\beta_{t}\\ &=\frac{\sqrt{\alpha_{t}}(1-\bar{\alpha}_{t-1})}{1-\bar{\alpha}_{t}}x^{\prime}_{t}+\frac{\sqrt{\bar{\alpha}_{t-1}}\beta_{t}}{1-\bar{\alpha}_{t}}x_{0}\\ &=\frac{\sqrt{\alpha_{t}}(1-\bar{\alpha}_{t-1})}{1-\bar{\alpha}_{t}}x^{\prime}_{t}+\frac{\sqrt{\bar{\alpha}_{t-1}}\beta_{t}}{1-\bar{\alpha}_{t}}\frac{1}{\sqrt{\bar{\alpha}_{t}}}(x^{\prime}_{t}-\sqrt{1-\bar{\alpha}_{t}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho}\boldsymbol{\epsilon}_{t})\\ &=\frac{1}{\sqrt{\alpha_{t}}}\Big{(}x^{\prime}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho}\boldsymbol{\epsilon}_{t}\Big{)}\end{split}$ (11)
According to Lemma 3.2, if we add Gaussian noise to the origin input image, which results in $\mathcal{N}(0,\rho^{2}I)$, then the distribution of generated images of the diffusion model has the same mean, but a variance scaled by $\frac{1}{\rho^{2}}$, where $\rho^{2}=2$ for clean samples.
Now we start analyzing the backdoor samples.
Definition G.5 (Backdoor Forward process).
Let $x_{0}^{b}\sim q(x_{b})$ denote a sample from target data distribution, $\delta$ denote a trigger, and $x_{T}^{b}\sim\mathcal{N}(\delta,I)$ denote the pure Gaussian noise attached by a trigger. Given the variance schedule $\{\beta_{t}\}_{t=1}^{T}$ in DDPM [22], define the forward process to diffuse $x_{0}^{b}$ to $x_{T}^{b}$ for backdoor samples:
$q(x_{t}^{b}|x_{t-1}^{b})=\mathcal{N}(x_{t}^{b};\sqrt{1-\beta_{t}}x_{t-1_{b}}+k_{t}\delta,\beta_{t}I),$ (12)
$q(x_{t}^{b}|x_{0}^{b})=\mathcal{N}(x_{t}^{b};\sqrt{\bar{\alpha}_{t}}x_{0}^{b}+\sqrt{1-\bar{\alpha}_{t}}\delta,(1-\bar{\alpha}_{t})I),$ (13)
where $k_{t}+\sqrt{\alpha_{t}}k_{t-1}+\sqrt{\alpha_{t}\alpha_{t-1}}k_{t-2}+...+\sqrt{\alpha_{t}...\alpha_{2}}k_{1}=\sqrt{1+\alpha_{t}}$.
By using a similar proof as for clean samples, we would easily derive a similar conclusion for backdoor samples: if we add Gaussian noise to the backdoor samples, the distribution of generated images of the diffusion model has the same mean, but $\frac{1}{\rho^{2}}$ variance to the original distribution.
In this paper, we only consider a simple case in the clean data distribution $q(x)$ follows some Gaussian distribution and leave more general cases in future works. Specifically, we consider that the clean data distribution $q(x)$ of the clean samples follow $\mathcal{N}(x_{c},\sigma_{c}I)$, while the backdoor samples follow $\mathcal{N}(x_{b},\sigma_{b}I)$.
For D2I attack, the target generated image is specific (e.g., Mickey Mouse), thus its variance is 0. For In-D2D and Out-D2D, the target images belong to the same class, hence their variance should be much lower than that of images generated from the clean model, as those are from various, completely different classes. Therefore, under the above case and Lemma 3.2, the distributions generated by clean and backdoor samples after noise addition are $\mathcal{N}(x_{c},\sigma_{c}\frac{1}{\rho^{2}}I)$ and $\mathcal{N}(x_{b},\sigma_{b}\frac{1}{\rho^{2}}I)$, respectively. In particular, we consider three attack methods. Hence, we have three kinds of $\sigma_{b}$. Let $\sigma_{D2I}$ denote the variance of D2I attack, $\sigma_{iD2D}$ denote the variance of In-D2I attack, $\sigma_{oD2I}$ denote the variance of Out-D2I attack defined in Appendix  B. We then have,
$0=\frac{1}{\rho^{2}}\sigma_{D2I}\ll\frac{1}{\rho^{2}}\sigma_{In-D2D},\frac{1}{\rho^{2}}\sigma_{Out-D2D}\ll\frac{1}{\rho^{2}}\sigma_{c}.$ (14)
Equation 14 shows the different variances of the images generated from the fixed diffusion model $f_{\theta}$ with noise added to the input in terms of clean and backdoor samples.
Under Lemma 3.1, which states that two points sample from a distribution of larger variance tend to have a larger distance than those from a distribution of lower variance, we have that after noise addition the backdoor inputs tend to generate similar images, while the clean inputs tend to generate diversified images. Take TrojDiff(In-D2D) attack on Cifar10 for example, $N$ is chosen as $32\times 32\times 3=3072$. Based on Lemma 3.1, we must have that $\|x_{1}-x_{2}\|_{2}\geq\|x_{3}-x_{4}\|_{2}$ under the condition of $\sigma_{In-D2D}<\frac{3071}{3072}\sigma_{c}$, where $x_{1},x_{2}\sim\mathcal{N}(x_{c},\sigma_{c}\frac{1}{\rho^{2}}I)$ and $x_{3},x_{4}\sim\mathcal{N}(x_{b},\sigma_{In-D2D}\frac{1}{\rho^{2}}I)$. However, according to Equation 14, this conditional is trivial to achieve, since $\frac{1}{\rho^{2}}\sigma_{c}\gg\frac{1}{\rho^{2}}\sigma_{In-D2D}$. This completes the proof.
Appendix H Proof of Lemma 3.2
Lemma H.1.
Let $f_{\theta}$ and $f_{\tilde{\theta}}$ be two well-trained diffusion models as defined in the Assumption G.4. Let $x^{\prime}_{T}$ follow $\mathcal{N}(0,\rho^{2}I)$. Let the clean data distribution be $q(x)\sim\mathcal{N}(x_{c},\sigma_{c}I)$. We then have:
$\hat{x_{0}}=f_{\theta}(x^{\prime}_{T})\sim\mathcal{N}(x_{c},\frac{\sigma_{c}}{\rho^{2}}I)$ (15)
Proof.
The output generated image from $f_{\tilde{\theta}}$ when input $x^{\prime}_{T}$ is given follows:
$\hat{\tilde{x_{0}}}=f_{\tilde{\theta}}(x^{\prime}_{T})\sim\mathcal{N}(x_{c},\sigma_{c}I),$ (16)
the Equation 16 is due to Assumption G.4. In particular, to obtain the generated image follows $q(x)$, the reverse process is defined as $q(x^{\prime}_{t-1}|x^{\prime}_{t})\sim\mathcal{N}(x^{\prime}_{t-1};\mu_{\tilde{\theta}}(x^{\prime}_{t},t),\Sigma_{\tilde{\theta}}(x^{\prime}_{t},t))$, where $\mu_{\tilde{\theta}}(x^{\prime}_{t},t)=\frac{1}{\sqrt{\alpha_{t}}}\Big{(}x^{\prime}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho}{\epsilon}_{t}\Big{)}$ and $\Sigma_{\tilde{\theta}}(x^{\prime}_{t},t)=\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\cdot\beta_{t}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho^{2}}$ (Equation 11). For the $f_{\theta}$, although the reverse process is also a gaussian distribution, $\mu_{\theta}(x_{t},t)=\frac{1}{\sqrt{\alpha_{t}}}\Big{(}x_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\boldsymbol{\epsilon}_{\theta}(x_{t},t)\Big{)}$, $\Sigma_{\theta}(x_{t},t)=\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\cdot\beta_{t}$ (Equation 6 and 7).
To obtain the generated image from $f_{\theta}$ when input $x^{\prime}_{T}$ is given, we substitute $x_{t}=x^{\prime}_{t}$ into the fixed $f_{\theta}$. We then have by Equation 9 that:
$\mu_{\theta}(x^{\prime}_{t},t)=\frac{1}{\sqrt{\alpha_{t}}}\Big{(}x^{\prime}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\boldsymbol{\epsilon}_{\theta}(x^{\prime}_{t}=\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho}\epsilon,t)\Big{)}$ (17)
$\Sigma_{\theta}(x^{\prime}_{t},t)=\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\cdot\beta_{t}$ (18)
Under the Assumption G.4, $\epsilon_{\theta}$ is able to accuaratly predict the noise added on the $\sqrt{\bar{\alpha}_{t}x_{0}}$ to obtain $x_{t}$, hence the prediction of $\epsilon_{\theta}$ in Equation 17 should be $\rho\epsilon_{t}$. We have by substituting $\rho\epsilon_{t}$ into Equation 17:
$\mu_{\theta}(x^{\prime}_{t},t)=\frac{1}{\sqrt{\alpha_{t}}}\Big{(}x^{\prime}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho}\epsilon_{t}\Big{)}$ (19)
By comparing Equation 19 and Equation 18 with Equation 11 , we found the mean of the reverse process is the same when inputting the $x^{\prime}_{T}$ to the $f_{\theta}$ and $f_{\tilde{\theta}}$, while the variance of $f_{\tilde{\theta}}$ is $\rho^{2}$ times larger than $f_{\theta}$. For simplicity, Let $a_{t}=\mu_{\theta}(x^{\prime}_{t},t)$ and $b_{t}=\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\cdot\beta_{t}$, we have: $q_{\theta}(x_{t-1}|x_{t})\sim\mathcal{N}(a_{t},b_{t}I)$ and $q_{\tilde{{\theta}}}(x_{t-1}|x_{t})\sim\mathcal{N}(a_{t},b_{t}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho^{2}}I)$. Hence, by the reparameterization trick, the variance of the generated $\hat{\tilde{x_{0}}}$ of $f_{\tilde{\theta}}$ is $\rho^{2}$ times greater than $f_{\theta}$. Without loss of generality, we use the Gaussian distribution to discribe the output distribution. Given the Assumption G.4, and $q(x)\sim\mathcal{N}(x_{c},\sigma_{c}I)$, $\hat{x_{0}}=f_{\theta}(x^{\prime}_{T})\sim\mathcal{N}(x_{c},\frac{\sigma_{c}}{\rho^{2}}I)$, which completes the proof. ∎
Appendix I Proof of Lemma 3.1
Lemma I.1 [10]).
Given that $x\sim\mathcal{N}(\mu,\sigma I)$, where $x$ is a $N$-dimensional vector. Then, we must have $\frac{N}{\sqrt{N+1}}\leq\sigma^{-1}\mathbb{E}(\|x\|_{2})\leq\sqrt{N}$, where $\mathbb{E}(X)$ denotes an expectation value of the random variable $X$.
Lemma I.2.
[…]
Proof.
Let $A:=\|x_{1}-x_{2}\|_{2}$, $B:=\|x_{3}-x_{4}\|_{2}$, $C:=B-A$, and $t$ be some arbitrary positive number. According to the Markov inequality, we have that
$\begin{split}Pr(C>t)\leq\frac{\mathbb{E}(C)}{t}=\frac{\mathbb{E}(B-A)}{t}=\frac{\mathbb{E}(B)-\mathbb{E}(A)}{t}=\frac{\mathbb{E}(\|x_{3}-x_{4}\|_{2})-\mathbb{E}(\|x_{1}-x_{2}\|_{2})}{t}\end{split}$ (20)
By basic calculations, it is obvious that $x_{1}-x_{2}\sim\mathcal{N}(0,\sqrt{2}\sigma_{1})$ and $x_{3}-x_{4}\sim\mathcal{N}(0,\sqrt{2}\sigma_{2})$. Then, according to Lemma I.1, we obtain that,
$Pr(C>t)\leq\sqrt{2}(\frac{\sigma_{2}\sqrt{N}}{t}-\frac{N\sigma_{1}}{t\sqrt{N+1}})\leq\sqrt{2}\frac{\sigma_{2}(N+1)-N\sigma_{1}}{t\sqrt{N+1}}$ (21)
If we have that $\sigma_{2}<\frac{N}{N+1}\sigma_{1}$, then the numerator will be always less than 0. This completes the proof. ∎

### v2

@@
-UFID: A Unified Framework for Input-level Backdoor Detection on Diffusion Models
+UFID: A Unified Framework for Black-box Input-level Backdoor Detection on Diffusion Models
 Abstract
-Diffusion Models are vulnerable to backdoor attacks, where malicious attackers inject backdoors by poisoning some parts of the training samples during the training stage.
-This poses a serious threat to the downstream users, who query the diffusion models through the API or directly download them from the internet.
-To mitigate the threat of backdoor attacks, there have been a plethora of investigations on backdoor detections.
-However, none of them designed a specialized backdoor detection method for diffusion models, rendering the area much under-explored.
-Moreover, these prior methods mainly focus on the traditional neural networks in the classification task, which cannot be adapted to the backdoor detections on the generative task easily.
-Additionally, most of the prior methods require white-box access to model weights and architectures, or the probability logits as additional information, which are not always practical.
-In this paper, we propose a Unified Framework for Input-level backdoor Detection (UFID) on the diffusion models, which is motivated by observations in the diffusion models and further validated with a theoretical causality analysis.
-Extensive experiments across different datasets on both conditional and unconditional diffusion models show that our method achieves a superb performance on detection effectiveness and run-time efficiency.
-The code is available at https://github.com/GuanZihan/official_UFID.
-1 Introduction
-Diffusion models [22, 40, 41, 38] have emerged as the new state-of-the-art family of generative models due to their superior performance [16] and wide applications across a variety of domains, ranging from computer vision [5, 8], natural language processing [2, 23, 28], and robust machine learning [7, 9].
-Despite their success, the training of diffusion models consumes a tremendous amount of time and computational resources.
-Consequently, it is a common practice to directly utilize third-party models via an API or directly download them from the internet.
-The setting is referred to as Model-as-a-Service (MaaS).
-Recently, it was found that diffusion models are vulnerable to backdoor attacks [11, 14, 15, 42, 20], where the malicious attackers poison some parts of the training samples with a predefined trigger pattern in the training stage.
-Consequently, the behavior of the diffusion models can be adversarially manipulated whenever the trigger pattern appears in the input query, while it remains normal with clean input queries.
-This vulnerability not only poses a serious threat to downstream users, such as inadvertently generating inappropriate images for children who query the model, but also poses risks to companies or artists with extensive copyright interests.
-For instance, it could manipulate the model to generate images that bypass copyright restrictions [43].
-Figure 1: Visualization of the input-level backdoor detection setting.
-To mitigate the threat of backdoor attacks, numerous investigations have been conducted focusing on backdoor detections [45, 31, 46].
-However, none of these studies specifically aim to develop a detection method for generative models, leaving this area significantly under-explored.
-Moreover, most of these methods are confined to traditional neural networks in classification tasks and are not readily adaptable to generative tasks due to two primary challenges: ❶ More Diverse Failures: Unlike merely generating a fixed target image (e.g., a Hello-Kitty image), backdoored diffusion models can be manipulated to produce a specific class of images (e.g., cat images), or even images with a specified abstract concept (e.g., erotic images).
+Diffusion models are vulnerable to backdoor attacks, where malicious attackers inject backdoors by poisoning certain training samples during the training stage.
+This poses a significant threat to real-world applications in the Model-as-a-Service (MaaS) scenario, where users query diffusion models through APIs or directly download them from the internet.
+To mitigate the threat of backdoor attacks under MaaS, black-box input-level backdoor detection has drawn recent interest, where defenders aim to build a firewall that filters out backdoor samples in the inference stage, with access only to input queries and the generated results from diffusion models.
+Despite some preliminary explorations on the traditional classification tasks, these methods cannot be directly applied to the generative tasks due to two major challenges: (1) more diverse failures and (2) a multi-modality attack surface.
+In this paper, we propose a black-box input-level backdoor detection framework on diffusion models, called UFID.
+Our defense is motivated by an insightful causal analysis: Backdoor attacks serve as the confounder, introducing a spurious path from input to target images, which remains consistent even when we perturb the input samples with Gaussian noise.
+We further validate the intuition with theoretical analysis.
+Extensive experiments across different datasets on both conditional and unconditional diffusion models show that our method achieves superb performance on detection effectiveness and run-time efficiency.
+Introduction
+Diffusion models (Ho, Jain, and Abbeel 2020; Song and Ermon 2020; Song et al.
+2020; Song, Meng, and Ermon 2020) have emerged as the new state-of-the-art family of generative models due to their superior performance (Dhariwal and Nichol 2021) and wide applications across a variety of domains, ranging from computer vision (Baranchuk et al.
+2021; Brempong et al.
+2022), natural language processing (Austin et al.
+2021; Hoogeboom et al.
+2021; Li et al.
+2022), and robust machine learning (Blau et al.
+2022; Carlini et al.
+2022).
+Despite their success, training diffusion models requires significant time and computational resources.
+Consequently, it is common practice to utilize third-party models via an API or to download them directly from the internet.
+This approach is known as Model-as-a-Service (MaaS).
+However, it has recently been found that diffusion models are vulnerable to backdoor attacks (Gu 2024; Huang et al.
+2024; Chen, Song, and Li 2023; Chou, Chen, and Ho 2023a, b; Struppek, Hintersdorf, and Kersting 2022; Guan et al.
+2023), where malicious attackers poison certain training samples with a predefined trigger pattern in the training stage.
+Consequently, the behavior of diffusion models can be adversarially manipulated whenever the trigger pattern appears in the input query, while it remains normal with clean input queries.
+This vulnerability poses a serious threat to real-world applications in the MaaS setting, e.g., the online third-party diffusion models may have been backdoored to generate inappropriate images for children or to generate images that bypass copyright restrictions when a specific trigger appears in the query  (Wang et al.
+2024a).
+More seriously, the traditional training-phase defense methods (Li et al.
+2021; Huang et al.
+2022; Gao et al.
+2023) cannot be deployed in the MaaS setting due to the defenders’ inaccessibility to the training pipeline and training data.
+To mitigate the above threat, black-box input-level backdoor detection has recently drawn great interest.
+In this scenario, input-level indicates that defenders aim to build a firewall-style detector in the inference stage to filter out and reject backdoored inputs while allowing clean inputs to generate predictions.
+Black-box means that defenders only have access to user queries and the generated results from the deployed models, without any prior information (e.g., model weights, architectures) assumed by the previous works (Ma et al.
+2022; Qiu et al.
+2021; Gao et al.
+2021, 2019).
+Prior methods for backdoor detection in image classification, e.g., (Guo et al.
+2023; Liu et al.
+2023; Hu et al.
+2024) cannot be adopted for generative tasks due to two major challenges: ❶ More Diverse Failures: Unlike merely generating a fixed target image (e.g., a Hello-Kitty image), backdoored diffusion models can be manipulated to produce a specific class of images (e.g., cat images), or even images with a specified abstract concept (e.g., erotic images).
 This implies that the target images are not necessarily unique but can vary as long as they belong to the designated target class.
-Consequently, the poisoning process differs significantly from simply inducing a poisoned model to predict a fixed target label in image classification tasks.
-This variability substantially complicates detection in generative tasks.
-❷ Multi-Modality Attack Surface: Diverging from traditional image classifiers, which involve a single modality, diffusion models (e.g., Stable Diffusion) are capable of supporting multiple modalities.
+This variability substantially complicates the detection in the generative task.
+❷ Multi-Modality Attack Surface: Unlike traditional image classifiers, which involve a single modality, diffusion models (e.g., Stable Diffusion) are capable of supporting multiple modalities.
 This diversity necessitates a unified framework for backdoor detection in diffusion models.
-Additionally, most of the prior methods require white-box access to the model weights and architecture, or the probability logits as additional information, which are not always practical in the MaaS setting.
-For the backdoor detection problem in the MaaS setting (i.e., black-box input-level backdoor detection [21, 30]; visualized in Figure 1), we assume no prior information about model weights and architectures, but that only the user query and the generated results from diffusion models are available.
-Intuitively, we aim to work as a firewall that receives users’ queries and determines whether the generated result can be returned to the users.
-In particular, we approve and forward the DNN’s prediction results for clean images while rejecting those for poisoned images.
-The challenges of the setting stem from two requirements: (1) efficiency: our method should not significantly influence the model’s inference speed, and (2) effectiveness: our method should effectively distinguish backdoor and clean samples.
-Therefore, how to detect backdoor samples in the MaaS setting for the diffusion models is far from a trivial question.
-To address the above challenges, we propose a Unified Framework for Input-level backdoor Detection (UFID) on diffusion models.
-Our method is motivated and underpinned by a causality analysis where the strong backdoor attack serves as the confounder, introducing a spurious path from input to target images.
-Furthermore, the spurious path embedded in the diffusion model remains consistent when we perturb the input samples with Gaussian noise.
-We further validate the causal analysis with a theoretical analysis.
-Driven by the analysis, we develop our final method for backdoor detection in diffusion models.
-Specifically, for conditional diffusion models, we gradually perturb the inputs with different random noises; for unconditional diffusion models, we perturb the text inputs by augmenting them with additional public textual prompts.
-Based on our previous analysis, a backdoored input tends to produce similar images regardless of perturbations, while a clean sample generates multiple diversified images.
-This contrast in image generation allows us to use the similarity of the generated images as a basis for detecting backdoor samples.
-Extensive experiments across different datasets on both conditional and unconditional diffusion models demonstrate that our method achieves superb performance on detection effectiveness and run-time efficiency.
-To sum up, our contributions in the work include: (1) First Unified Backdoor Detection Framework for Diffusion Models.
-To the best of our knowledge, our work is the first unified framework for detecting backdoor samples in diffusion models; (2) Novel Causality Analysis and Theoretical Analysis.
-We are the first paper to apply causality analysis for analyzing backdoor attacks on generative tasks; Besides, we also provide a theoretical analysis to validate the effectiveness of our method; (3) SOTA performance.
-Extensive results on various datasets empirically show that our method achieves outstanding performance in terms of detection effectiveness and runtime efficiency.
-2 Preliminaries
-2.1 Diffusion Models
-Without loss of generality, diffusion models are composed of two components: (1) Diffusion Process: a data distribution $q(x)$ is diffused to a target distribution $r(x)$ within T timestamps.
-(2) Training Process: a diffusion model with parameter $\theta$ is trained to align with the reversed diffusion process, i.e., $p_{\theta}(x_{i-1}|x_{i})=\mathcal{N}(x_{i-1};\mu_{\theta}(x_{i}),\sigma_{\theta}(x_{i}))=q(x_{i-1}|x_{i})$.
-In this section, we introduce the basic version DDPM [22].
-DDPM.
+An overall comparison of the problem setting is in the Appendix.
+To address the above challenges, our intuition is motivated by a causal analysis: backdoor attacks serve as the confounder, introducing a spurious path from input to target images.
+The spurious path embedded in the diffusion model remains consistent even when we perturb input samples with Gaussian noise.
+Therefore, the backdoor generations remain consistent after minor perturbation, while the clean generations alter significantly even with a small perturbation.
+We further validate the causal analysis rigorously, showing that after perturbation, the difference between the diversity of clean generations and the diversity of backdoor generations is lower bounded (Corollary 3).
+Driven by the analysis, we develop a Unified Framework for black-box Input-level backdoor Detection (UFID) on diffusion models.
+Specifically, UFID examines each input sample by calculating its graph density score, which measures the similarity within the generated batch after perturbing the given input sample.
+The higher the graph density score, the more likely the input sample is backdoored.
+Compared to the existing method TERD (Mo et al.
+2024), UFID is designed for a black-box setting, requiring no access to model weights or structures.
+Moreover, the performance gap between UFID and TERD is also satisfactory, with a maximum difference of 8% in precision, 7% in Recall, and 1% in AUC.
+UFID is also generalized to the scenario of conditional diffusion models, where the input can be of various modalities.
+To strengthen UFID’s resistance to diversity-intensive backdoor attacks, we also design strategies to integrate supplementary correspondence information into the UFID framework.
+In contrast, Shield (Wang et al.
+2024b) operates in a white-box setting but achieves similar performance as UFID.
+To sum up, our contributions in the work include: (1) First Unified Black-box Backdoor Detection Framework for Diffusion Models.
+To the best of our knowledge, our work is the first unified black-box framework for detecting input-level backdoor samples in diffusion models; (2) Novel Causality Analysis.
+We apply causality analysis for analyzing backdoor attacks on generative tasks; Besides, theoretical analysis also sheds light on our intuitions and validates the effectiveness of our method; (3) Promising Performance.
+Extensive results show that our detection method achieves an average of nearly 100% AUC on the unconditional models, and 90% AUC on the conditional models with an acceptable inference overhead.
+[^1] An extended version of this work with appendices and further details is available at: https://arxiv.org/abs/2404.01101
+Related Works
+Backdoor Attacks and Defenses on Diffusion Models.
+Recently, a lot of works have investigated the security vulnerabilities of diffusion models by launching backdoor attacks on diffusion models.
+From a high-level idea, malicious backdoor attackers aim to inject a special behavior in the diffusion process such that, once the predefined trigger pattern appears on the input, the special behaviors will be activated.
+To achieve this goal,  (Chen, Song, and Li 2023) proposed to add an additional backdoor injection task on the training stage and maliciously alter the sampling procedure with a correction term.
+(Chou, Chen, and Ho 2023a) proposed a novel attacking strategy by only modifying the training loss function.
+(An et al.
+2023) focused on launching attacks to text-to-image tasks, by injecting backdoors into the pre-trained text encoder.
+(Huang et al.
+2024) proposed to apply personalization techniques to efficiently inject a malicious concept into the diffusion models.
+(Chou, Chen, and Ho 2023b) proposed a unified framework that covers all the popular schemes of diffusion models, including conditional and unconditional diffusion models.
+(Zhai et al.
+2023) proposed a novel backdoor attack that can generate backdoor images as diversified as clean images.
+Backdoor defenses on diffusion models are highly under-explored.
+To the best of our knowledge, only four papers (An et al.
+2023; Mo et al.
+2024; Sui et al.
+2024; Wang et al.
+2024b) investigated backdoor defenses on diffusion models.
+However, Elijah (An et al.
+2023) aims to detect whether a given model is backdoored, while our tasks focus on filtering backdoor samples for diffusion models in the inference stage, which are fundamentally different.
+TERD (Mo et al.
+2024) builds a unified framework for safeguarding diffusion models, which can handle tasks such as trigger inversion, input detection, and model detection.
+However, the proposed technique only works for unconditional diffusion models and requires white-box access to the weights and structures of the diffusion models.
+DisDet (Sui et al.
+2024) also only focuses on unconditional diffusion models.
+It conducts backdoor detection by checking whether the given input follows a Gaussian distribution or not.
+However, the defense method can be easily circumvented by using an imperceptible trigger pattern.
+(Wang et al.
+2024b) proposed to detect backdoors on the text-to-image models based on a novel ”assimilation phenomenon”.
+Preliminaries
+Diffusion Models.
+Without loss of generality, diffusion models contain two parts: (1) Diffusion Process: a data distribution $q(x)$ is diffused to a target distribution $r(x)$ within $T$ timestamps.
+(2) Training Process: A diffusion model $\epsilon_{\theta}$ with parameter $\theta$ is trained to align with the reversed diffusion process, i.e., $p_{\theta}(x_{i-1}|x_{i})=\mathcal{N}(x_{i-1};\mu_{\theta}(x_{i}),\sigma_{\theta}(x_{i}))=q(x_{i-1}|x_{i})$.
+DDPM is one of the most basic diffusion models (Ho, Jain, and Abbeel 2020).
 DDPM assumes the target distribution $r(x)=\mathcal{N}(0,I)$ and the diffusion process $q(x_{i}|x_{t-1})=\mathcal{N}(x_{t};\sqrt{1-\beta_{i}}x_{i-1},\beta_{i}I)$, where the $\{\beta_{i}\}_{i=1}^{T}$ is a pre-defined variance schedule that controls the step sizes.
@@
 In the inference stage, DDPM generates images by sampling from the Gaussian distribution $\mathcal{N}(0,I)$ from time $i=T$ to $i=0$ with the generative process $p_{\theta}(x_{i-1}|x_{i})=\mathcal{N}(x_{i-1};\mu_{\theta}(x_{i}),\sigma_{\theta}(x_{i}))$, where $\mu_{\theta}(x_{i})=\frac{1}{\alpha_{i}}(x_{i}-\frac{1-\alpha_{i}}{1-\bar{\alpha_{i}}\epsilon_{\theta}(x_{i},i)})$ and $\sigma_{\theta}(x_{i})=\frac{(1-\bar{\alpha}_{i-1})\beta_{i}}{1-\bar{\alpha}_{i}}$.
-2.2 Backdoor Attacks on Diffusion Models
-Different from launching backdoor attacks to the traditional models (e.g., classifiers [19, 12]), which could be achieved by poisoning training dataset, injecting backdoors into the diffusion models is much more complicated.
-A typical backdoor attack pipeline [11, 14, 15] on diffusion models consists of three steps: (1) the attackers first need to mathematically define the forward backdoor diffusion process, i.e., $x_{0}^{b}\rightarrow x^{b}_{T}$, where the $x_{0}^{b}$ denotes the target image and the $x_{T}^{b}$ denotes the trigger image; (2) then the attackers train the diffusion models to align with the backdoored reversed process; (3) in the inference stage, the diffusion models can be prompted to generate the target images when the input contains the trigger pattern, but behave normally when the input is clean (e.g., pure gaussian noise for the DDPM model).
-2.3 Threat Model
-We adopt a similar setting as in  [21, 30], which is named "black-box input-level backdoor detection".
-Specifically, two parties are considered in the setting: Attacker and Defender.
-We present a detailed description of each party as follows.
-Attacker.
-We assume a strong attacker that is capable of training a backdoored model and uploading the backdoored model to the public models market, such as cloud platforms, GitHub, and Hugging Face.
-The downstream users, such as software engineers, can then be authorized to deploy the model in the local end device.
-In the inference stage, the malicious attacker can activate the backdoor by querying the deployed model with an additional trigger in the input.
-Defender.
-The defender aims to conduct efficient and effective input-level black-box backdoor detection on the user side.
-Specifically, efficiency requires that the detection process does not significantly impact the response time of user queries, while effectiveness requires the detection process to distinguish backdoor samples and clean samples with a high accuracy rate.
-The defender is assumed to only have access to the deployed model’s outputs (e.g., the generated image from the diffusion model), without any prior information such as model weight and embedding.
-Problem 1 (Input-level Backdoor Detection).
-For a given diffusion model represented as $f_{\theta}(\cdot)$, the defender’s objective of the defender is to develop a detector, denoted by $\mathcal{A}(\cdot)$.
-This detector aims to output $\mathcal{A}(\bm{x}_{backdoor})=1$ for poisoned inputs, and $\mathcal{A}(\bm{x}_{clean})=0$ for clean inputs.
-The challenge of Problem 1 arises from three factors:
-  1.
-More Diverse Failures.
+Backdoor Attacks on Diffusion Models.
+Different from launching backdoor attacks on the traditional models (e.g., classifiers (Gu, Dolan-Gavitt, and Garg 2017; Chen et al.
+2017)), which could be achieved by poisoning training dataset, injecting backdoors into the diffusion models is much more complicated.
+A typical backdoor attack pipeline (Chen, Song, and Li 2023; Chou, Chen, and Ho 2023a, b) on diffusion models consists of three steps: (1) the attackers first need to mathematically define the forward backdoor diffusion process, i.e., $x_{0}^{b}\rightarrow x^{b}_{T}$, where the $x_{0}^{b}$ denotes the target image and the $x_{T}^{b}$ denotes the trigger image; (2) then the attackers train the diffusion models to align with the backdoored reversed process; (3) in the inference stage, the diffusion models can be prompted to generate target images when the input contains the trigger pattern, but behave normally when the input is clean (e.g., pure Gaussian noise for the DDPM model).
+Threat Model.
+We adopt a similar threat model as in (Guo et al.
+2023; Gao et al.
+2021; Liu et al.
+2023).
+Specifically, the defender is assumed to only have access to user queries (e.g., prompt) and the generated results from diffusion models.
+The defender aims to conduct efficient and effective black-box backdoor detection, where efficiency requires that the detection process does not significantly impact the response time of user queries, while effectiveness requires the detection process to distinguish backdoor samples and clean samples with a high accuracy rate.
+The challenge of this threat model arises from three factors: (1) More Diverse Failures.
 Backdoored diffusion models can be triggered to generate specific classes of images (e.g., cat images), or even images with a specified abstract concept (e.g., erotic images), extending beyond fixed target labels in traditional classification tasks.
-  2. Multi-Modality Attack Surface. Unlike traditional image classifiers that involve only one modality, diffusion models (e.g., Stable Diffusion) can support a variety of modalities.
-  3.
-Limited Information & Efficiency Guarantee.
-The detector $\mathcal{A}(\cdot)$ has access only to the query images and the prediction labels returned by the DNN, and is expected to minimally impact the inference efficiency.
-3 Methodology
+(2) Multi-Modality Attack Surface.
+Unlike traditional image classifiers that involve only one modality, diffusion models (e.g., Stable Diffusion) can support a variety of modalities.
+(3) Limited Information.
+The detection method only has access to the query images and the prediction labels returned by the diffusion model.
+Overview of the UFID
+Intuition: Backdoor Attacks under a Causal Lens
+Figure 1: Causal graph of clean and backdoored generation.
 Figure 2: Pipeline of our unified framework for backdoor detection on diffusion models.
-3.1 Backdoor Attacks from a Causal View
 To develop a backdoor detection algorithm for generative models, we first need to address a fundamental question: What distinguishes clean generation from backdoored generation, and how this distinction can be utilized in designing an effective detection algorithm?
-To this end, we propose to leverage causal inference to reveal the distinct mechanisms underlying clean and backdoored generation processes.
-Specifically, we present a causal graph that illustrates the comparison of the two processes in Figure 3.
-A causal graph is a directed acyclic graph that illustrates the causal relationships among variables, where each node represents a variable.
-An edge from a node $A$ to a node $B$ implies that the variable $A$ is the cause of the variable $B$ (denoted as $A\rightarrow B$).
-For simplicity, this figure only includes the unconditional diffusion model.
-However, this model can be easily extended to the conditional diffusion model by substituting the input image ($X_{T}$) with the input text ($T$).
-Figure 3: Causal graph of clean and backdoored generation.
+Motivated by the great potential of causal inference in deep learning, we propose to leverage causal inference as a new perspective to understand the distinct mechanisms underlying clean and backdoored generation processes.
+Specifically, we construct causal graphs to illustrate the comparison between the two processes, as shown in Figure 1.
+A causal graph is a directed acyclic graph that illustrates the causal relationships among variables, where each node represents a variable and each edge represents a causal relationship.
+For simplicity, this figure only illustrates the unconditional diffusion model.
+However, the causal graph can be easily extended to the conditional diffusion model by substituting the input noise ($X_{T}$) with the input text ($T$).
 Clean Generation.
-As depicted in Figure 3(a), the generated image ($x_{0}^{c}$) is dependent on the input noise image $x_{T}^{c}\sim\mathcal{N}(0,I)$.
+As depicted in Figure 1(a), the generated image ($x_{0}^{c}$) is dependent on the input noise $x_{T}^{c}\sim\mathcal{N}(0,I)$.
 This relationship is termed the causal path, denoted as $X_{T}^{c}\rightarrow X_{0}^{c}$.
-Consequently, adding Gaussian noise $\epsilon\sim\mathcal{N}(0,I)$ to $x_{T}^{c}$ results in a new input ${x^{\prime}}_{T}^{c}=x_{T}^{c}+\epsilon=\mathcal{N}({x^{\prime}}_{T}^{c};0,2I)$, leading to a different generated image ${x^{\prime}}_{0}^{c}$, as shown in Figure 3(c).
+Consequently, adding a small Gaussian noise $\epsilon\sim\mathcal{N}(0,I)$ with a small weight $\alpha$ to $x_{T}^{c}$ results in a new input ${x^{\prime}}_{T}^{c}=x_{T}^{c}+\alpha\cdot\epsilon=\mathcal{N}({x^{\prime}}_{T}^{c};0,(1+\alpha)I)$, leading to a different generated image ${x^{\prime}}_{0}^{c}$, as shown in Figure 1(c).
 Backdoored Generation.
-As shown in Figure 3(b), a backdoor attack $A$ modifies an image $x_{T}$ by injecting a trigger $\delta$ and changing the image generation process towards the target image $x_{0}^{b}$, denoted as $X_{T}^{b}\leftarrow A\rightarrow X_{0}^{b}$, where $x_{T}^{b}=\delta+x_{T}$.
-This introduces a spurious path from $X_{T}^{b}$ to $X_{0}^{b}$, which lies outside the direct causal path $X_{T}^{b}\rightarrow X_{0}^{b}$, thereby establishing and strengthening erroneous correlations between the modified input noise and the target image.
-Consequently, generations from poisoned noise images are primarily influenced by this spurious path [17, 47, 29], while the direct causal path $X_{T}^{b}\rightarrow X_{0}^{b}$ plays a minor role, represented by a gray dotted line in Figure 3 (b).
-When an additional Gaussian noise $\epsilon\sim\mathcal{N}(0,I)$ is added to the backdoored noise image $x_{T}^{b}$, the new backdoored input becomes a Gaussian noise ${x^{\prime}}_{T}^{b}\sim\mathcal{N}(\delta,2I)$, which is a combination of a new noise image $x_{T}+\epsilon$ and the trigger pattern $\delta$.
-Since the trigger pattern remains in the new backdoored input, the spurious path continues to dominate the generation process.
-As a result, the generated image will remain unchanged.
+As shown in Figure 1(b), a backdoor attack $A$ modifies an image $x_{T}$ by injecting a trigger $\delta$ and changing the image generation process towards the target image $x_{0}^{b}$, denoted as $X_{T}^{b}\leftarrow A\rightarrow X_{0}^{b}$, where $x_{T}^{b}=\delta+x_{T}^{c}$.
+This introduces a spurious path from $X_{T}^{b}$ to $X_{0}^{b}$, which lies outside the direct causal path $X_{T}^{b}\rightarrow X_{0}^{b}$, thereby establishing and strengthening the erroneous correlation between the modified input noise and the target image.
+Consequently, generations of poisoned noise images are primarily influenced by this spurious path (Du et al.
+2021; Zhang et al.
+2023; Li et al.
+2021), while the direct causal path $X_{T}^{b}\rightarrow X_{0}^{b}$ plays a minor role, represented by a gray dotted line in Figure 1 (b).
+When an additional Gaussian noise $\epsilon\sim\mathcal{N}(0,I)$ is added to the backdoored noise image $x_{T}^{b}$ with a small weight $\alpha$, the new backdoored input becomes ${x^{\prime}}_{T}^{b}\sim\mathcal{N}(\delta,(1+\alpha)I)$, which is a combination of a new noise image $x_{T}^{c}+\alpha\cdot\epsilon$ and the trigger pattern $\delta$.
+It can be interpreted as poisoning a new image $x_{T}^{c}+\alpha\cdot\epsilon$ with the trigger $\delta$.
+Hence, the generated image is still affected by the attack and lies within the domain of target images.
+In addition, the magnitude of the perturbation is controlled by a small weight $\alpha$ (e.g., 0.01), ensuring that the trigger pattern remains in the new input without being disrupted.
 In summary, for clean generation, a small perturbation significantly alters the output.
@@
 Consequently, minor perturbations of backdoor samples do not lead to substantial changes in the diffusion model’s generation results.
-The following theorems validate our insights from causal analysis.
-Lemma 3.1.
+The following theorems also validate our insights from causal analysis.
+Lemma 1.
+Let $f_{\theta}$ and $f_{\tilde{\theta}}$ be two well-trained diffusion models as defined in Assumption 11 in the Appendix.
+Let input noise $x^{\prime}_{T}$ follow $\mathcal{N}(0,\rho^{2}I)$.
+Let $\hat{x_{0}}$ be the generated image for $x^{\prime}_{T}$ and the generated distribution for clean input $x_{T}^{c}$ be $q(x)\sim\mathcal{N}(x_{c},\sigma_{c}I)$.
+We then have:
+$\hat{x_{0}}=f_{\theta}(x^{\prime}_{T})\sim\mathcal{N}(x_{c},\frac{\sigma_{c}}{\rho^{2}}I).$ (1)
+The lemma implies that if we modify the variance of the input noise by multiplying it by $\rho^{2}$ (adding small Gaussian noise), the variance of the generated image is reduced to $\frac{1}{\rho^{2}}$ of the original.
+Theorem 2.
+Suppose the output domain of the diffusion model $f_{\theta}$ is Gaussian.
+Let $\mathcal{N}(x_{c},\sigma_{c})$ and $\mathcal{N}(x_{b},\sigma_{b})$ denote the distribution of clean generations and backdoor generations respectively.
+Let $N$ denote the image size.
+We assume that $\sigma_{c}\geq\sigma_{b}+\rho^{2}$.
+Given clean input noise $x_{T}^{c}\sim\mathcal{N}(0,I)$, backdoor input noise $x_{T}^{b}=x_{T}^{c}+\delta\sim\mathcal{N}(\delta,I)$, and Lemma 1 if we perturb the clean input noises $x_{T}$ and backdoor input noise $x_{T}^{b}$ with some $\epsilon\sim\mathcal{N}(0,I)$ simultaneously, then for the resulted clean generations $\mathcal{N}(x^{\prime}_{c},\sigma^{\prime}_{c})$ and the backdoor generations $\mathcal{N}(x^{\prime}_{b},\sigma^{\prime}_{b})$, we have that $\sigma^{\prime}_{c}-\sigma^{\prime}_{b}\geq 1$.
+Corollary 3.
 […]
-The Lemma 3.1 states that two points sampled from a distribution with larger variance tend to be more diversified than those sampled from a distribution with lower variance.
-A detailed proof is provided in Appendix H.
-Lemma 3.2.
-Let $f_{\theta}$ and $f_{\tilde{\theta}}$ be two well-trained diffusion models as defined in the Assumption G.4.
-Let $x^{\prime}_{T}$ follow $\mathcal{N}(0,\rho^{2}I)$.
-Let the clean data distribution be $q(x)\sim\mathcal{N}(x_{c},\sigma_{c}I)$.
-We then have:
-$\hat{x_{0}}=f_{\theta}(x^{\prime}_{T})\sim\mathcal{N}(x_{c},\frac{\sigma_{c}}{\rho^{2}}I)$ (1)
-This Lemma 3.2 indicates that after adding Gaussian noise to the input, the mean of generated image keeps unchanged, while the variance is scaled by $\frac{1}{\rho^{2}}$.
-In our paper, $\rho^{2}=2$.
-A detailed proof is provided in Appendix I.
-Theorem 3.3.
-Given a diffusion model $f_{\theta}$, clean input noise $x_{T}\sim\mathcal{N}(0,I)$, and backdoor input noise $x_{T}^{b}=x_{T}+\delta\sim\mathcal{N}(\delta,I)$, if we perturb the clean input noise and backdoor input noise with some $\epsilon\sim\mathcal{N}(0,I)$ simultaneously, then the generated images from clean input noise will be much more diversified than that from backdoor input noise.
-Proof.
-The proof is under Lemma 3.1 and  3.2. We provide detailed proof in Appendix G.
-3.2 Scenario 1: Unconditional Diffusion Models
+$\mathbb{E}(\|x_{1}-x_{2}\|_{2}-\|x_{3}-x_{4}\|)\geq\frac{N(\sigma_{c}-\sigma_{b})-\sigma_{b}}{\sqrt{N+1}}>0.$ (2)
+Theorem 2 and Corollary 3 imply that, after adding noise, the expected difference between the distance of generated clean images $\|x_{1}-x_{2}\|_{2}$ and the distance of the generated backdoor images $\|x_{3}-x_{4}\|_{2}$ is significantly larger than 0.
+This further sheds light on the intuition that the diversity of clean generations is significantly greater than backdoor generations after adding perturbations.
+Their proofs are provided in the Appendix.
+Scenario 1: Unconditional Diffusion Models
 […]
-$y_{i}^{j}={f_{\theta}(x_{i}^{j})},\forall x_{i}^{j}\in\mathbb{I},$ (2)
-We further let $y_{i}=\{y_{i}^{j}|1\leq j\leq|\mathbb{M}|+1\}$ denote the generated batch. We can then determine whether the input query $x_{i}$ is a backdoored sample or not by inspecting $y_{i}$.
-3.3 Scenario 2: Conditional Diffusion Models
+$y_{i}^{j}={f_{\theta}(x_{i}^{j})},\forall x_{i}^{j}\in\mathbb{I}.$ (3)
+Let $y_{i}=\{y_{i}^{j}|1\leq j\leq|\mathbb{M}|+1\}$ denote the generated batch. We can then determine whether the input query $x_{i}$ is a backdoored sample by inspecting the diversity of $y_{i}$.
+Scenario 2: Conditional Diffusion Models
 Conditional diffusion models accept input from other modalities to guide the generation of user-intended images.
-For instance, in stable diffusion [36], textual input is used to query the model.
+For instance, in stable diffusion (Rombach et al.
+2021), textual input is used to query the model.
 However, it is evident that the detection approach employed for unconditional diffusion models cannot be directly applied to conditional diffusion models due to the inability to introduce Gaussian noise to discrete textual input.
@@
 We leverage variations in output diversity to distinguish between clean and backdoor samples.
-To further accentuate this diversity gap and enhance distinguishability in conditional setting, we propose appending the input text $x_{i}$ with a random phrase $ph_{j}$ selected from a diverse phrase pool containing completely distinct phrases, such as "Iron Man" and "Kitchen Dish Washer," denoted as =$\mathbb{N}=\{ph_{1},ph_{2},...ph_{|\mathbb{N}|}\}$, where $|\mathbb{M}|\ll|\mathbb{N}|$.
-For each input $x_{i}$, we repeat this process $|\mathbb{M}|$ times, resulting in augmented an input batch $\mathbb{I}=\{x_{i}\}\cup\{x_{i}^{j}|x_{i}^{j}=x_{i}\oplus ph_{j},\forall 1\leq j\leq|\mathbb{M}|\}$, where $\oplus$ denotes the string appending operator.
+To further enlarge this diversity gap and enhance distinguishability in conditional setting, we propose appending the input text $x_{i}$ with a random phrase $ph_{j}$ selected from a diverse phrase pool containing completely distinct phrases, such as ”Iron Man” and ”Kitchen Dish Washer,” denoted as =$\mathbb{N}=\{ph_{1},ph_{2},...ph_{|\mathbb{N}|}\}$.
+For each input $x_{i}$, we repeat this process $|\mathbb{M}|$ times, where $|\mathbb{M}|\ll|\mathbb{N}|$.
+This results in an augmented input batch $\mathbb{I}=\{x_{i}\}\cup\{x_{i}^{j}|x_{i}^{j}=x_{i}\oplus ph_{j},\forall 1\leq j\leq|\mathbb{M}|\}$, where $\oplus$ denotes the string appending operator.
 Similarly, we can generate an image batch $y_{i}$ by querying the target stable diffusion model using Equation 3.
-3.4 A Unified Framework for Backdoor Detection
-As previously discussed, the diversity of images generated by a specific input $i$ can be leveraged to detect backdoors.
+A Unified Framework for Backdoor Detection
+As previously discussed, the diversity of images generated by a specific input $i$ can be used to detect backdoors.
 To quantify this diversity, we employ a two-step process that involves calculating both pairwise and overall similarities within the generated batch.
@@
 Pairwise Similarity Calculation.
-Initially, we calculate semantic embeddings of generated images through a pre-trained image encoder (e.g., ViT-ImageNet [44] and CLIP [35]), denoted as $f_{\mathcal{E}}(\cdot)$.
+Initially, we calculate semantic embeddings of generated images through a pre-trained image encoder (e.g., ViT-ImageNet (Wu et al.
+2020) and CLIP (Radford et al.
+2021)), denoted as $f_{\mathcal{E}}(\cdot)$.
 Subsequently, we calculate the local similarity for each pair of images in the generated batch using cosine similarity, represented by $S_{c}(\cdot,\cdot)$.
 Graph Density Calculation.
-Following this, we construct a weighted graph and compute its graph density to represent the overall similarity of all images within graph.
+Following this, we construct a weighted graph and compute its graph density to represent the overall similarity of all images within the graph.
 Specifically, let $G_{i}=(V_{i},\mathcal{E}_{i})$ represent the similarity graph for input sample $x_{i}$, where $|V_{i}|=|\mathbb{M}|$ constitutes the set of vertices (symbolizing the generated images) and $\mathcal{E}_{i}$ is the set of edges.
@@
 In this similarity graph, the similarity between two generated images is interpreted as the distance within the graph.
-Next, we introduce graph density [4], as a novel metric for evaluating the overall similarity of the generated batch:
-Definition 3.4.
+Next, we introduce graph density (Balakrishnan and Ranganathan 2012), as a novel metric for evaluating the overall similarity of the generated batch:
+Definition 4.
 The graph density $DS(G_{i})$ of the weighted similarity graph is defined as:
@@
 If $DS(G_{i})$ is greater than the threshold $\tau$, then it is determined as a backdoor sample, otherwise, it is a clean sample and the originally generated image shall be returned to the users.
-The total pipeline of our method is visualized in Figure 2 and the final detection algorithm is presented in Algorithm 1.
-4 Experimental Results
-4.1 Experimental Settings
+The total pipeline of our method is visualized in Figure 2 and the final detection algorithm is in Algorithm 1.
+Remark 5 (Reliance on Pre-trained Encoders).
+Relying on a pre-trained encoder $f_{\mathcal{E}}(\cdot)$ might not always be feasible in practice.
+To relax this assumption, we also explore using a model-free metric structural similarity index measure (SSIM) to calculate the pairwise similarity between images.
+We report the evaluation results in the Appendix.
+Remark 6 (Applicability of UFID).
+The effectiveness of UFID is based on a practical assumption that backdoor generations are more similar than clean generations, which has been implicitly made in the previous backdoor attacks work, e.g., the backdoor generations share a similar style (Struppek, Hintersdorf, and Kersting 2022; Huang, Guo, and Juefei-Xu 2023), object (Chou, Chen, and Ho 2023b, a), or semantic concept (Chen, Song, and Li 2023).
+It is also observed that when the target images of backdoor attacks are as diversified as the clean images (Zhai et al.
+2023), the performance of UFID will be limited.
+However, we could incorporate supplementary information to enhance the UFID.
+Detailed evaluations are provided in the Appendix.
+Experiments
+Experimental Settings
 Attack Baselines.
-To the best of our knowledge, the existing backdoor attacks on diffusion models include two unconditional-DM-based backdoor attacks: TrojDiff [11] and BadDiffusion [14], and two conditional-DM-based backdoor attacks: Rickrolling [42] and Villandiffusion [15].
-We consider all four backdoor attacks as our attack baselines.
-It is also noted that TrojDiff supports three types of attack objectives.
-Detailed descriptions of the four attack methods are provided in Appendix B.
+To the best of our knowledge, the existing backdoor attacks on diffusion models include two unconditional-DM-based backdoor attacks: TrojDiff (Chen, Song, and Li 2023) and BadDiffusion (Chou, Chen, and Ho 2023a), and three conditional-DM-based backdoor attacks: Rickrolling (Rick) (Struppek, Hintersdorf, and Kersting 2022), Villandiffusion (VillanDiff) (Chou, Chen, and Ho 2023b), and Personalization (Personal) (Huang, Guo, and Juefei-Xu 2023).
+We consider all five backdoor attacks as our attack baselines.
+It is also noted that TrojDiff, Rickrolling, and Personalization all support diversity-preserving backdoor attacks, where the attackers’ target images are diversified.
+Detailed descriptions of them are provided in the Appendix.
 Defense Baselines.
-To the best of our knowledge, there are yet methods for detecting backdoor samples on diffusion models.
-Moreover, existing methods are built especially for classification models and therefore not applicable to the generative models.
+We compared our method with the existing well-established defense method TERD (Mo et al.
+2024) on the unconditional diffusion models and Shield (Wang et al.
+2024b) on the conditional diffusion models.
+It is noted that both TERD and Shield require additional white-box access to the model weights and structures, which are not always feasible in our MaaS setting.
 Models and Datasets.
 Different backdoor attacks are built based on different backbone models and samplers.
-To facilitate evaluation, TrojDiff [11] and BadDiffusion [15] are evaluated on DDPM [22], while VillanDiffusion [15] and Rickrolling [42] are evaluated on Stable Diffusion [36].
-For the training datasets, we choose Cifar10 [26] and CelebA [32] for TrojDiff and BadDiffusion, and choose CelebA-HQ-Dialog [25] and Pokemon [34] for VillanDiffusion and Rickrolling.
+To facilitate evaluation, TrojDiff and BadDiffusion are evaluated on DDPM, while VillanDiffusion, Rickrolling, and Personalization are evaluated on Stable Diffusion v1.4 (Rombach et al.
+2021).
+For the training datasets, we choose CIFAR10 (Krizhevsky, Nair, and Hinton ) and CelebA (Liu et al.
+2015) for TrojDiff and BadDiffusion, and choose CelebA-D (Jiang et al.
+2021) and Pokemon (Pinkney 2022) for VillanDiffusion, Rickrolling, and Personalization.
 Metrics.
 Following the prior works on backdoor detection, we adopt three popular metrics for evaluating the effectiveness of our detection method: Precision (P), Recall (R), and Area under the Receiver Operating Characteristic (AUC).
-Implementation Details
-All the models are well-trained with the default hyper-parameters reported in the original papers so that they show a good performance in generating both clean images and backdoor images.
-Following the previous works [27, 21], we then evaluate our detection method with a positive (i.e., attacked) and a negative (i.e., clean) dataset.
-For evaluations against unconditional diffusion models, we randomly generate 1000 Gaussian noises as the clean queries (negative) and construct backdoor samples (positive) accordingly by blending the trigger pattern with the Gaussian noises.
-For evaluations against conditional diffusion models, we split the whole dataset into 90% train and 10% test following [15].
-Then, we use the textual caption in the test subset as the clean queries (negative) and construct backdoor queries (positive) accordingly.
-The threshold value $\tau$ is determined by a small clean held-out validation dataset, where detailed descriptions are provided in Appendix C.
-The pre-trained encoder is set as CLIP-ViT-B32 [35], the magnitude size is set as 4, the poisoning rate is set as 10%, and the number of validation datasets is set as 20, by default.
-All the hyperparameters are evaluated in the ablation studies.
-4.2 Main Results
-Effectiveness
+Implementation Details.
+All the models are well-trained with the default hyper-parameters reported in the original papers.
+Following the previous works (Lee et al.
+2018; Guo et al.
+2023), we evaluate our detection method with a positive (i.e., attacked) and a negative (i.e., clean) dataset.
+Due to space limits, the details for constructing the two datasets and the default hyper-parameters are provided in the Appendix.
+Main Results
+Effectiveness.
+UFID(black-box) TERD(white-box)
+Dataset Attacks P R AUC P R AUC
+Cifar10 TrojDiff(D2I) 0.95 0.94 1.00 1.00 1.00 1.00
+TrojDiff(In) 0.93 0.93 0.98 1.00 1.00 1.00
+TrojDiff(Out) 0.93 0.92 1.00 1.00 1.00 1.00
+BadDiffusion 0.93 0.95 1.00 1.00 1.00 1.00
+Average 0.93 0.94 1.00 1.00 1.00 1.00
+CelebaA TrojDiff(D2I) 0.93 0.92 1.00 1.00 1.00 1.00
+TrojDiff(In) 0.90 0.89 0.96 1.00 1.00 1.00
+TrojDiff(Out) 0.91 0.92 0.98 1.00 1.00 1.00
+BadDiffusion 0.97 0.95 1.00 1.00 1.00 1.00
+Average 0.93 0.92 0.99 1.00 1.00 1.00
 Table 1: Performance of the proposed detection method against backdoor attacks on unconditional diffusion models.
-Dataset Backdoor Attacks Precision Recall AUC
-Cifar10 TrojanDiff(D2I) 0.95 0.94 1.00
-TrojanDiff(In-D2D) 0.93 0.93 0.98
-TrojanDiff(Out-D2D) 0.93 0.92 1.00
-BadDiffusion 0.93 0.95 1.00
-Average 0.93 0.94 1.00
-CelebaA TrojanDiff(D2I) 0.93 0.92 1.00
-TrojanDiff(In-D2D) 0.90 0.89 0.96
-TrojanDiff(Out-D2D) 0.91 0.92 0.98
-BadDiffusion 0.97 0.95 1.00
-Average 0.93 0.92 0.99
+UFID(black-box) Shield(white-box)1
+Dataset Attacks P R AUC P R AUC
+CelebaA-D VillanDiff 0.92 0.95 0.96 0.80 0.95 -
+Rick(TPA) 0.87 0.84 0.90 0.96 0.85 -
+Rick(TAA) 0.82 0.81 0.87 - - -
+Personal 0.76 0.72 0.82 - - -
+Average 0.85 0.84 0.89 0.88 0.90
+Pokemon VillanDiff 0.91 0.93 0.94 - - -
+Rick(TPA) 0.83 0.85 0.91 - - -
+Rick(TAA) 0.80 0.81 0.87 - - -
+Personal 0.73 0.77 0.81 - - -
+Average 0.82 0.85 0.89 - - -
+  • 1
+    Due to unavailable codes, we use the reported values in the original paper directly.
 Table 2: Performance of the proposed detection method against backdoor attacks on conditional diffusion models.
-Dataset Backdoor Attacks Precision Recall AUC
-CelebaA-HQ-Dialog VillanDiffusion 0.92 0.95 0.96
-Rickrolling 0.87 0.84 0.90
-Average 0.90 0.90 0.93
-Pokemon Caption VillanDiffusion 0.91 0.93 0.94
-Rickrolling 0.83 0.85 0.91
-Average 0.87 0.89 0.93
-Table 2 presents the performance of our detection method against backdoor attacks on unconditional diffusion models, and Table 2 presents the performance on conditional diffusion models.
-As shown, the AUC values for different backdoor attack methods on all the evaluated datasets are over 0.9, suggesting that our method can effectively distinguish backdoor and clean samples.
-Efficiency
-Figure 4: Average inference speed of the proposed detection method against TrojDiff(D2I) on the Cifar10 dataset.
-Figure 4 illustrates the efficiency of our detection method against TrojDiff(D2I) on the Cifar10 dataset.
+Table 1 presents the performance of our detection method against backdoor attacks on unconditional diffusion models, while Table 2 presents the performance on conditional diffusion models.
+As shown, the AUC values for different backdoor attack methods on all the evaluated datasets are over 0.8, suggesting that our method can effectively distinguish backdoor and clean samples.
+Compared to the baseline methods, UFID shows a comparable performance with a slight drop.
+However, the decrease is reasonable as the two baselines are white-box methods, which require additional access to the model weights and structures.
+The similar performance drop has been observed in the previous studies (Guo, Li, and Liu 2021; Guo et al.
+2023).
+Figure 3: Average inference speed against TrojDiff(D2I) on the Cifar10.
+Figure 4: Average inference speed against VillanDiffusion on the Pokemon dataset.
+Efficiency.
+Figure 3 illustrates the efficiency of our detection method against TrojDiff(D2I) on CIFAR10 (32$\times$32$\times$ 3) dataset.
 We query the diffusion models with 320 samples with a batch size of 64 and record the average inference speed, defined as the average time consumption for a sample.
@@
 Due to the increased number of query samples, UFID inevitably results in a lower inference speed than that of the vanilla mode.
-However, the increased time consumption is within expectations and will not significantly influence user experience in practice.
-4.3 Ablation Studies
-In this section, we discuss how the hyper-parameters influence the effectiveness of the detection method.
+However, the increased time consumption is within expectations.
+Figure 4 presents the efficiency of the UFID against VillanDiffusion on the Pokemon (512$\times$512$\times$3) dataset.
+Due to the acceleration techniques in the modern sampler (Song, Meng, and Ermon 2020; Lu et al.
+2022), stable diffusion models are usually denoised with only a small number of steps (e.g., 25) to generate high-resolution images.
+Therefore, we could observe that the UFID only slightly increases the inference time.
+To further investigate how the selected inference steps influence the efficiency, we evaluate the inference speed with different numbers of inference steps.
+The results show that when setting the inference step from 10 to 75, the inference overhead is acceptable.
+Ablation Studies
+In this section, we discuss how the hyper-parameters influence the effectiveness of the UFID.
 The Influence of Different Pre-trained Encoders.
-Table 3: Performance of our detection method with different pre-trained encoders.
-Encoder $\rightarrow$ ViT-ImageNet CLIP DINO V2
-ViT-B ViT-L RN50 RN50x64 ViT-B ViT-L ViT-S ViT-B ViT-L
-Precision 0.94 0.95 0.89 0.85 0.91 0.80 0.81 0.85 0.80
-Recall 0.93 0.95 0.88 0.81 0.91 0.79 0.70 0.78 0.65
-AUC 0.98 0.99 0.88 0.95 0.97 0.88 0.99 1.00 0.99
-The pre-trained encoder is a crucial component in our detection method.
-To evaluate its impact on the effectiveness of our detection method, we test the performance of UFID when integrated with different pre-trained encoders in Table 3.
-For the space limit, we only present the results against TrojDiff(In-D2D) on the Cifar10 dataset in the main manuscript.
-More experiments are in the Appendix D.
-As shown, UFID works well when integrated with different pre-trained encoders.
+The pre-trained encoder is important in our detection method.
+To evaluate its impact on the effectiveness of our detection method, we test the performance of UFID when integrated with different pre-trained encoders.
+For the space limit, we present the results in the Table 3 - Table 8 in the Appendix.
+According to the tables, UFID works well when integrated with different pre-trained encoders.
 In particular, CLIP encoders show consistently good performance across different datasets, due to their strong generalization ability from the pre-training stage.
@@
 Figure 5: Performance with different sizes of magnitude set.
-Figure 6: Performance with different amounts of available samples.
-Figure 6 investigates the impact of the magnitude size on the running-time efficiency and effectiveness, where the X-axis denotes the magnitude size, the left y-axis denotes the performance values, and the right y-axis denotes the average inference speed.
-The first four backdoor attacks are evaluated on the Cifar10 dataset and the remaining two backdoor attacks are evaluated on the Pokemon dataset.
+Figure 5 investigates the impact of the magnitude size on the running-time efficiency and effectiveness, where the left y-axis denotes the performance values, and the right y-axis denotes the average inference speed.
+Due to the space limit, we only present partial results in the main manuscript, while the remaining parts can be found in the Appendix.
 As the figure shows, our detection method achieves a stably satisfactory performance against all backdoor attacks when the size is over four.
@@
 The Influence of Available Validation Dataset.
-Figure 6 presents the impact of the available validation dataset on the threshold values and the performance, where the X-axis values denote the number of available validation samples and the y-axis denotes the performance values.
-The first four backdoor attacks are evaluated on the Cifar10 dataset and the remaining two backdoor attacks are evaluated on the Pokemon dataset.
-As the figure suggests, with more validation samples available, the threshold values gradually increase, and the performances tend to become more stable.
+Figure 16 and Figure 17 in the Appendix presents the impact of the available validation dataset on the performance, where the X-axis values denote the number of available validation samples and the y-axis denotes the performance values.
+As the figure suggests, with more validation samples available, the performances tend to become more stable.
 However, it is also noted that if the number of validation samples becomes exceedingly large, there is a slight drop in performance.
 A possible explanation for this phenomenon is that more validation samples also introduce more noisy information, leading to an unexpected threshold value.
+The Influence of Image Size.
+Figure 20 in the Appendix investigates whether UFID’s performance will be influenced when handling high-resolution images.
+Specifically, we evaluate UFID against TrojDiff and BadDiffusion on an augmented CIFAR10 dataset, where the image size is manually scaled to 64, 128, and 256.
+We could see that the image size does not have any influence on the performance, demonstrating UFID’s potential to handle high-resolution images.
 The Influence of Different Poisoning Rate.
-Figure 8 explores whether the performance of UFID is sensitive to the backdoor poisoning rate.
+Figure 15 in the Appendix explores whether the performance of UFID is sensitive to the backdoor poisoning rate.
 We evaluate the UFID under poisoning rate from 0.05 to 0.30.
-The results reveal that our method can generally provide a satisfactory performance irrelevant to the poisoning rate.
-Moreover, with an increase in the poisoning rate, the performance becomes more stable.
-Figure 7: Performance with different poisoning rates.
-Figure 8: Evaluation of UFID against adaptive attacks.
+The results reveal that our method can perform satisfactorily under different poisoning rates.
+Moreover, with the increase of the poisoning rate, the performance becomes more stable.
+Discussions
+Visualizations of Similarity Graphs.
+To better understand how UFID helps to detect backdoor samples, we visualize the similarity graphs in Figure 6.
+Due to the space limit, we only present similarity graphs against the TrojDiff(In-D2D) on CIFAR10 in the main manuscript.
+More qualitative examples are presented in the appendix.
+Each node in the similarity graph denotes the generated images of the query batch, while each edge denotes the cosine similarity scores of the embedding of any two images.
+As shown, the similarity scores for the clean query batch are significantly lower than those for the backdoor query batch, validating our intuitions for backdoor detection.
+Visualizations of Scores Distributions.
+Figure 18 and Figure 19 in the Appendix present the distributions of the graph density $\mathcal{S}_{i}$ for backdoor samples and clean samples respectively.
+As shown, the distribution of backdoor samples tends to be more clustered in a narrow range, while that of clean samples tends to be spread out.
+Moreover, there is a distinct gap between the two distributions, suggesting that UFID can effectively distinguish backdoor and clean samples.
 Resilient against Adaptive Backdoor Attacks.
+Figure 6: Similarity graphs against TrojDiff (In-D2D) attack on CIFAR10 dataset.
 We evaluate our detection method against an adaptive attacker who already has prior information about our detection method.
@@
 In this way, the attacker achieves a more stealthy backdoor attack by sacrificing the attack success rate.
-We further define the ratio between the number of clean images to the backdoor samples in this target domain as the "blending ratio".
+We further define the ratio between the number of clean images to the backdoor samples in this target domain as the ”blending ratio”.
 We evaluate the UFID against TrojDiff(D2I) and employ mean square error (MSE) between the generated backdoor images and the target image (e.g., Mickey Mouse) as the attack success rate.
-Figure 8 presents the performance of the UFID under different blending ratios.
-As shown, the performance of the UFID exhibits a slight decrease when the blending ratio rises.
+Figure 21 in the Appendix presents the performance of the UFID under different blending ratios.
+As shown, the UFID’s performance gradually decreases when the blending ratio rises.
 However, the average MSE across generated backdoor samples abruptly exceeds 0.15, which suggests the failure in injecting backdoors.
 The right-hand side also provides some samples in the CIFAR10 dataset, where we notice that images with an MSE of 0.15 to the target image are already completely different from the target image.
-4.4 Discussions
-Visualizations of Similarity Graphs.
-Figure 9: Similarity graphs against TrojDiff (In-D2D) attack on Cifar10 dataset.
-To better understand how UFID helps to detect backdoor samples, we visualize the similarity graphs in Figure 9.
-Due to the space limit, we only present similarity graphs against the TrojDiff In-D2D attack on the Cifar10 dataset in the main manuscript.
-More qualitative examples are presented in the appendix E.
-Each node in the similarity graph denotes the generated images of the query batch, while each edge denotes the cosine similarity scores of the embedding of any two images.
-As shown, the similarity scores for the clean query batch are significantly lower than those for the backdoor query batch, validating our intuitions for backdoor detections.
-Visualizations of Scores Distributions.
-Figure 17 and Figure 18 present the distributions of the calculated scores $\mathcal{S}_{i}$ for backdoor samples and clean samples respectively.
-As shown, the distribution of backdoor samples tends to be more clustered in a narrow range, while that of clean samples tends to be spread out.
-Moreover, there is a distinct gap between the two distributions, suggesting that our method can effectively distinguish backdoor and clean samples.
-5 Related Works
-Diffusion Models.
-Diffusion models have become a new state-of-the-art family of generative models in image synthesis by showing strong sample quality and diversity in image synthesis [22, 38].
-Since then, diffusion models have been widely applied to different types of tasks, such as image generation [39, 41], image super-resolution [37, 6], and image editing [3, 13].
-Moreover, researchers also found that the representations learned by diffusion models can also be used in other discriminative tasks such as segmentation [18] and anomaly detection [33].
-Backdoor Attacks and Defenses on Diffusion Models.
-Recently, inspired by the great success of diffusion models, a lot of works investigate the security vulnerabilities of diffusion models by launching backdoor attacks on diffusion models.
-From a high-level idea, malicious backdoor attackers aim to inject a special behavior in the diffusion process such that, once the predefined trigger pattern appears on the input, the special behaviors will be activated.
-To achieve this goal,  [11] proposed to add an additional backdoor injection task on the training stage and maliciously alter the sampling procedure with a correction term.
-[14] proposed a novel attacking strategy by only modifying the training loss function.
-[1] focused on launching attacks to text-to-image tasks, by injecting backdoors into the pre-trained text encoder.
-[15] proposed a unified framework that covers all the popular schemes of diffusion models, including conditional diffusion models and unconditional diffusion models.
-However, backdoor defenses on diffusion models are highly under-explored.
-To the best of our knowledge, only  [1] investigated backdoor defenses on diffusion models.
-However, their work aims to detect whether a given model is backdoored, while our tasks focus on filtering backdoor samples for diffusion models in the inference stage, which are fundamentally different.
-6 Conclusion and Future Directions
+Conclusion and Future Directions
 In this paper, we propose a simple unified framework for backdoor detection on diffusion models under the MaaS setting.
-Our framework is first motivated by a strict causality analysis on image generation and further validated by theoretical analysis.
+Our framework is first motivated by a causality analysis on image generation and further validated by theoretical analysis.
 Motivated by the analysis, we design a unified method for distinguishing backdoor and clean samples for both conditional and unconditional diffusion models.
@@
 Despite the great success, there are still many directions to be explored in the future.
-Firstly, our method relies heavily on a pre-trained encoder, which is not always practical.
-It is promising to design some model-free metrics to detect backdoor samples on diffusion models.
-Secondly, extending our unified framework to the newly emerging diffusion models is also of great interest.
-Impact Statment
-Diffusion Models have been widely adopted for generating high-quality images and videos.
-Therefore, inspecting the security of diffusion models is of great significance in practice.
-In this paper, we propose a simple unified framework that effectively detects backdoor samples for the diffusion models under a strict but practical scenario of Moel-as-a-Service (MaaS).
-As described in the threat model, our method is proposed from the perspective of a defender.
-Therefore, this paper has no ethical issues and will not introduce any additional security risks to diffusion models.
-However, it is noted that our method is only used for filtering backdoored testing samples but they do not reduce the intrinsic backdoor vulnerability of the deployed diffusion models.
-We will further improve our method in future works.
+For example, UFID still requires a small amount of validation samples to determine the threshold.
+Can we relax the assumption?
+Acknowledgments
+The work is supported in part by the U.S.
+Office of Naval Research Award under Grant Number N00014-24-1-2668 and the National Science Foundation under Grants IIS-2316306, CNS-2330215, CCF-1918656, CNS-2317193, IIS-2331315.
 Appendix A Pseudo code of the UFID Detection Algorithm
@@
 Algorithm 1 Backdoor Detection on Diffusion Models
-  Input: User Input $x_{i}$; Target Diffusion Model $f_{\theta}$; Detection threshold $\tau$.
+  Input: User Input $x_{i}$; Target Diffusion Model $f_{\theta}$; Detection threshold $\tau$; small weight $\alpha=0.01$.
   if unconditional model then
-     $\mathbb{I}=\{x_{i}\}\cup\{x_{i}^{j}|x_{i}^{j}=x_{i}+\epsilon_{j},\forall 1\leq j\leq|\mathbb{M}|\}$
+     $\mathbb{I}=\{x_{i}\}\cup\{x_{i}^{j}|x_{i}^{j}=x_{i}+\alpha\cdot\epsilon_{j},\forall 1\leq j\leq|\mathbb{M}|\}$
   else if conditional model then
@@
   end if
-  for $j=1$ to $|\mathcal{M}|$ do
+  for $j=1$ to $|\mathcal{M}|+1$ do
      $y_{i}^{j}={f_{\theta}(x_{i}^{j})},\forall x_{i}^{j}\in\mathbb{I}$
@@
 Appendix B More Details about Attack Baselines
-TrojDiff [11].
+TrojDiff (Chen, Song, and Li 2023).
 We implement TrojDiff following the public code [^2] https://github.com/chenweixin107/TrojDiff on GitHub.
@@
 Figure 7: Examples of backdoor samples from TrojDiff.
-BadDiffusion [14].
+BadDiffusion (Chou, Chen, and Ho 2023a).
 We implement BadDiffusion following the public code [^3] https://github.com/FrankCCCCC/baddiffusion_code/tree/master on GitHub.
@@
 Figure 8: Examples of backdoor samples from BadDiffusion.
-VillanDiffusion [15].
+VillanDiffusion (Chou, Chen, and Ho 2023b).
 We implement VillanDiffusion following the public code [^4] https://github.com/IBM/VillanDiffusion/tree/main on GitHub.
@@
 Specifically, the backdoor attacks is conducted over a pre-trained stable diffusion model [^5] https://huggingface.co/CompVis/stable-diffusion-v1-4/tree/main, so as to make the model generates target images once the caption trigger appears.
-We use "mignneko" as the caption trigger, and the Cat image used in the original paper as the target image, since these configurations are shown to perform well on different datasets in the original paper.
+We use ”mignneko” as the caption trigger, and the Cat image used in the original paper as the target image, since these configurations are shown to perform well on different datasets in the original paper.
 We give an illustration of VillanDiffusion in the Figure 9.
 Figure 9: Examples of backdoor samples from VillanDiffusion.
-Rickrolling [42].
+Rickrolling (Struppek, Hintersdorf, and Kersting 2022).
 We implement Rickrolling following the public code [^6] https://github.com/LukasStruppek/Rickrolling-the-Artist on GitHub.
@@
 It uses the Cyrillic o as the trigger and replaces $o$ in the original text to construct backdoor samples.
-The target text is chosen as "a drawing of a bird with blue eyes".
+It supports two attack modes: Target Prompt Attacks (TPA) and Target Attribute Attacks (TAA), respectively.
+For the TPA, the target text is chosen as ”a drawing of a bird with blue eyes”, while for the TAA, the target text is chosen as ”black and white photo”.
 We give a illustration of Rickrolling in the Figure 10.
 Figure 10: Examples of backdoor samples from Rickrolling.
-Appendix C More Details about How to Choose $\tau$.
-Suppose we are given $n$ clean validation samples: $x_{1}$, $x_{2}$, …, $x_{n}$, then we take them as a batch and query the diffusion model as described in 2.
+Personalization (Huang et al. 2024)
+We implemented Personalization following the instructions [^7] https://github.com/Huang-yihao/Personalization-based__\__backdoor on Gihutb.
+Our implementation follows the pipeline in textual inversion, but we construct the training dataset with mismatched text-image pairs.
+Specifically, we use $\langle dog\rangle$ as the trigger and the cat toy images [^8] https://huggingface.co/datasets/valhalla/images as the target images.
+The Personalization is originally implemented to directly attack pre-trained Stable Diffusion models.
+To align with our experimental settings, we fine-tune the Stable Diffusion models on the Pokemon/CelebA-D dataset before conducting backdoor injection.
+Appendix C More Implementation Details
+Overall Implementation
+All the models are well-trained with the default hyper-parameters reported in the original papers so that they show a good performance in generating both clean images and backdoor images.
+Following the previous works (Lee et al.
+2018; Guo et al.
+2023), we then evaluate our detection method with a positive (i.e., attacked) and a negative (i.e., clean) dataset.
+For evaluations against unconditional diffusion models, we randomly generate 1000 Gaussian noises as the clean queries (negative) and construct backdoor samples (positive) accordingly by blending the trigger pattern with the Gaussian noises.
+For evaluations against conditional diffusion models, we split the whole dataset into 90% train and 10% test following (Chou, Chen, and Ho 2023b).
+Then, we use the textual caption in the test subset as the clean queries (negative) and construct backdoor queries (positive) accordingly.
+Following a practical assumption in backdoor detection (Guo, Li, and Liu 2021), the threshold value $\tau$ is determined by a small clean hold-out validation dataset, where detailed descriptions are provided in the next section.
+The pre-trained encoder is set as CLIP-ViT-B32 (Radford et al.
+2021), the size of magnitude set is chosen as 4, the poisoning rate is set as 10%, and the number of validation datasets is set as 20, by default.
+All the hyperparameters are evaluated in the ablation studies.
+More Details about How to Choose $\tau$.
+Suppose we are given $n$ clean validation samples: $x_{1}$, $x_{2}$, …, $x_{n}$, then we take them as a batch and query the diffusion model as described in 3.
 In this way, a similarity graph $G$ can be constructed on this batch, with each edge denoting the similarity between any two generated images.
@@
 The maximal average value is used as the threshold $\tau$.
+Computational Resources
+We conduct all the experiments on a server with 4×4$\times$4 × 80GB NVIDIA A100s.
 Appendix D More Ablation Studies on Pre-trained Encoders
 In this section, we record performances of our detection method UFID against different backdoor attacks when integrated with different pre-trained encoders, where Table 4 is for TrojDiff(Out-D2D), Table 5 is for TrojDiff(D2I), Table 6 is for BadDiffusion, Table 7 is for VillanDiffusion, and Table 8 is for Rickrolling.
-Table 4: Performance of our detection method against TrojDiff(Out-D2D) on the Cifar10 dataset with different pre-trained encoders.
+Encoder $\rightarrow$ ViT-ImageNet CLIP DINO V2
+ViT-B ViT-L RN50 RN50x64 ViT-B ViT-L ViT-S ViT-B ViT-L
+Precision 0.94 0.95 0.89 0.85 0.91 0.80 0.81 0.85 0.80
+Recall 0.93 0.95 0.88 0.81 0.91 0.79 0.70 0.78 0.65
+AUC 0.98 0.99 0.88 0.95 0.97 0.88 0.99 1.00 0.99
+Table 3: Performance of our detection method with different pre-trained encoders.
 Encoder $\rightarrow$ ViT-ImageNet CLIP DINO V2
@@
 AUC 1.00 1.00 1.00 0.99 1.00 1.00 1.00 1.00 1.00
-Table 5: Performance of our detection method against TrojDiff(D2I) on the Cifar10 dataset with different pre-trained encoders.
+Table 4: Performance of our detection method against TrojDiff(Out-D2D) on CIFAR10 dataset with different pre-trained encoders.
 Encoder $\rightarrow$ ViT-ImageNet CLIP DINO V2
@@
 AUC 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00
-Table 6: Performance of our detection method against BadDiffusion on the Cifar10 dataset with different pre-trained encoders.
+Table 5: Performance of our detection method against TrojDiff(D2I) on CIFAR10 dataset with different pre-trained encoders.
 Encoder $\rightarrow$ ViT-ImageNet CLIP DINO V2
@@
 AUC 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00 1.00
-Table 7: Performance of our detection method against VillanDiffusion on the Pokemon dataset with different pre-trained encoders.
+Table 6: Performance of our detection method against BadDiffusion on CIFAR10 dataset with different pre-trained encoders.
 Encoder $\rightarrow$ ViT-ImageNet CLIP DINO V2
@@
 AUC 0.64 0.68 0.97 0.94 0.94 0.90 0.91 0.90 0.92
-Table 8: Performance of our detection method against Rickrolling on the Pokemon dataset with different pre-trained encoders.
+Table 7: Performance of our detection method against VillanDiffusion on the Pokemon dataset with different pre-trained encoders.
 Encoder $\rightarrow$ ViT-ImageNet CLIP DINO V2
@@
 AUC 0.66 0.68 0.94 0.90 0.91 0.89 0.86 0.87 0.84
+Table 8: Performance of our detection method against Rickrolling on the Pokemon dataset with different pre-trained encoders.
 Appendix E More Details about Similarity Graphs
 We provide additional qualitative examples of similarity graphs in Figure 11, Figure 12 and Figure 13.
-Specifically, Figure 14 presents similarity graphs for backdoor attacks on the Cifar10 dataset, where the leftmost image represents a similarity graph for a clean query.
+Specifically, Figure 11 presents similarity graphs for backdoor attacks on CIFAR10 dataset, where the leftmost image represents a similarity graph for a clean query.
 Moving from left to right, we present qualitative examples of similarity graphs for backdoor queries under TrojDiff(D2I), TrojDiff(Out-D2D), and BadDiffusion, respectively.
 Moreover, Figure 12 and Figure 13 present similarity graphs for VillanDiffusion and Rickrolling backdoor attacks, where the left image represents a similarity graph for a clean query, and the right image is a similarity graph for a backdoor query.
-Figure 14: Similarity graphs generated for backdoor attacks on the Cifar10 dataset.
+Figure 11: Similarity graphs generated for backdoor attacks on CIFAR10 dataset.
 The leftmost image represents a similarity graph for a clean query.
@@
 Figure 13: Similarity graphs generated for Rickrolling backdoor attacks. The left image represents a similarity graph for a clean query, and the right image is a similarity graph for a backdoor query.
-Appendix F More Details about Score Distributions
+Appendix F More Ablation Studies on Magnitude Set
+Figure 14 presents the impact of the size of magnitude set on the performance against Personalization.
+Figure 14: Performance with different magnitude size.
+Appendix G More Ablation Studies on Poisoning Rates
+Figure 15 presents the impact of the poisoning rate on the performance of the UFID against different types of backdoor attacks.
+Note that Personalization injects backdoors with only 3-5 samples, without any definitions of ”poisoning rate”.
+Figure 15: Performance with different poisoning rates.
+Appendix H More Ablation Studies on Available Samples
+Figure 16 and Figure 17 present the impact of the available validation dataset on the performance of UFID.
+Figure 16: Performance with different amounts of available samples.
+Figure 17: Performance with different amounts of available samples.
+Appendix I More Details about Score Distributions
 In Figure 18, we provide distributions of graph density scores $DS(G)$ for both clean and backdoor samples on CIFAR10 dataset against TrojDiff(D2I), TrojDiff(Out-D2D), TrojDiff(In-D2D), and BadDiffusion attack, where the red bars denote the scores for clean samples, and the blue bars denote the scores for backdoor samples.
@@
 For all of the distributions, we can notice there exists an obvious gap between the score distributions for backdoor samples and those for clean samples, suggesting that our detection method can effectively distinguish backdoor and clean samples.
-Figure 17: Distributions of detection scores for backdoor samples and clean samples against unconditional diffusion models.
+Appendix J Ablation Studies on Image Size
+In Figure 20, we evaluate UFID against TrojDiff and BadDiffusion on an augmented CIFAR10 dataset, where the image size in CIFAR10 dataset is manually scaled to 64, 128, and 256.
+We could see that the image size does not have any influence on the performance, demonstrating UFID’s potential to handle high-resolution images.
+Figure 18: Distributions of detection scores for backdoor and clean samples on unconditional diffusion models.
 Figure 19: Distributions of detection scores for backdoor samples and clean samples against conditional diffusion models.
-Appendix G Proof of Theorem 3.3
-Theorem G.1.
-Given diffusion model $f_{\theta}$, clean input noise $x_{T}\sim\mathcal{N}(0,I)$, and backdoor input noise $x_{T}^{b}=x_{T}+\delta\sim\mathcal{N}(\delta,I)$, if we perturb the clean input noise and backdoor input noise with some $\epsilon\sim\mathcal{N}(0,I)$ simultaneously, then the generated images from clean input noise will be more diversified than that from backdoor input noise.
+Figure 20: Performance with different Image Size.
+Appendix K Adaptive Attacks
+Figure 21 shows the performance of UFID against adaptive attacks.
+Figure 21: Evaluation of UFID against adaptive attacks.
+Appendix L Evaluations with Diversity-intensive Backdoor Attacks.
+BadT2I (Zhai et al.
+2023) is a novel diversity-intensive backdoor attack method.
+BadT2I contains three modes: Pixel-backdoor, Object-backdoor, and Style-backdoor, respectively.
+For simplicity, we start our analysis based on the most difficult one, i.e., object-backdoor here.
+The analysis can be easily applied to the other two modes.
+Detailed descriptions of the other two modes can be found in the original paper.
+[…]
+The proposed UFID pipeline is not able to effectively distinguish backdoor generations from clean generations since the backdoor generations (i.e., bike images) can be as diversified as the clean generations.
+For example, in Figure 22, the graph density scores for the clean generations and backdoor generations are hardly distinguished.
+Despite the great challenge, we found that additional correspondence information between the input prompt and the generations could be integrated into the existing UFID pipeline without violating the black-box assumptions in the threat model.
+[…]
+$\texttt{Corre}(x_{i},y_{i})=-\langle\texttt{CLIP}(y_{i}),\texttt{CLIP}(x_{i})\rangle$ (4)
+Figure 22: Graph density score distributions of backdoor samples and clean samples.
+Figure 23: Detection score distributions of backdoor samples and clean samples.
+Then the final detection score can be the sum of the Corre score and the graph density score.
+It is noted that the graph density score here is manually scaled with a weight $(|\mathbb{M}|-1)$ to match the scale of the Corre score, where $|\mathbb{M}|$ is the size of the generated batch.
+We conduct experiments on three types of BadT2I: pixel-backdoor, object-backdoor, and style-backdoor.
+For each type of BadT2I, we ask ChatGPT to generate 100 random prompts according to their default specifications.
+The backdoor samples are constructed by concatenating the trigger with each prompt.
+The following Table 9 shows the AUC values of our detection performance.
+Pixel-backdoor Object-backdoor Style-backdoor
+UFID (w/ Corre) 0.90 0.94 0.86
+Table 9: AUC values of our enhanced UFID method on three types of BadT2I backdoor attacks.
+Appendix M Social Impact Statment
+Diffusion Models have been widely adopted for generating high-quality images and videos.
+Therefore, inspecting the security of diffusion models is of great significance in practice.
+In this paper, we propose a simple unified framework that effectively detects backdoor samples for the diffusion models under a strict but practical scenario of Moel-as-a-Service (MaaS).
+As described in the threat model, our method is proposed from the perspective of a defender.
+Therefore, this paper has no ethical issues and will not introduce any additional security risks to diffusion models.
+However, it is noted that our method is only used for filtering backdoored testing samples but they do not reduce the intrinsic backdoor vulnerability of the deployed diffusion models.
+We will further improve our method in future works.
+Appendix N Experiments about Model-free Similarity Metric
+To relax the assumption of using a pre-trained encoder for calculating image similarities, we explore model-free metrics like SSIM for image similarity.
+Table 10 reports additional experiments on CIFAR10 with TrojDiff, demonstrating that UFID is effective with SSIM as the similarity metric.
+Table 10: Effectiveness of UFID on Cifar10 dataset with SSIM.
+Attack P R AUC
+TrojDiff(D2I) 0.95 0.96 1.00
+TrojDiff(Out) 0.93 0.91 0.98
+TrojDiff(In) 0.73 0.76 0.84
+Figure 24: Comparison of the problem settings.
+Appendix O Proof of Lemma 1
+Lemma 7.
+Let $f_{\theta}$ and $f_{\tilde{\theta}}$ be two well-trained diffusion models as defined in the Assumption 11.
+Let input noise $x^{\prime}_{T}$ follow $\mathcal{N}(0,\rho^{2}I)$.
+Let $\hat{x_{0}}$ be the generated image for $x^{\prime}_{T}$.Let the clean data distribution be $q(x)\sim\mathcal{N}(x_{c},\sigma_{c}I)$.
+We then have:
+$\hat{x_{0}}=f_{\theta}(x^{\prime}_{T})\sim\mathcal{N}(x_{c},\frac{\sigma_{c}}{\rho^{2}}I)$ (5)
+Proof.
+The output generated image from $f_{\tilde{\theta}}$ when input $x^{\prime}_{T}$ is given follows:
+$\hat{\tilde{x_{0}}}=f_{\tilde{\theta}}(x^{\prime}_{T})\sim\mathcal{N}(x_{c},\sigma_{c}I),$ (6)
+the Equation 6 is due to Assumption 11.
+In particular, to obtain the generated image follows $q(x)$, the reverse process is defined as $q(x^{\prime}_{t-1}|x^{\prime}_{t})\sim\mathcal{N}(x^{\prime}_{t-1};\mu_{\tilde{\theta}}(x^{\prime}_{t},t),\Sigma_{\tilde{\theta}}(x^{\prime}_{t},t))$, where $\mu_{\tilde{\theta}}(x^{\prime}_{t},t)=\frac{1}{\sqrt{\alpha_{t}}}\Big{(}x^{\prime}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho}{\epsilon}_{t}\Big{)}$ and $\Sigma_{\tilde{\theta}}(x^{\prime}_{t},t)=\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\cdot\beta_{t}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho^{2}}$ (Equation 8).
+For the $f_{\theta}$, although the reverse process is also a gaussian distribution, $\mu_{\theta}(x_{t},t)=\frac{1}{\sqrt{\alpha_{t}}}\Big{(}x_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\boldsymbol{\epsilon}_{\theta}(x_{t},t)\Big{)}$, $\Sigma_{\theta}(x_{t},t)=\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\cdot\beta_{t}$ (Equation 13 and 14).
+To obtain the generated image from $f_{\theta}$ when input $x^{\prime}_{T}$ is given, we substitute $x_{t}=x^{\prime}_{t}$ into the fixed $f_{\theta}$. We then have by Equation 16 that:
+$\mu_{\theta}(x^{\prime}_{t},t)=\frac{1}{\sqrt{\alpha_{t}}}\Big{(}x^{\prime}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\boldsymbol{\epsilon}_{\theta}(x^{\prime}_{t}=\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho}\epsilon,t)\Big{)}$ (7)
+$\Sigma_{\theta}(x^{\prime}_{t},t)=\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\cdot\beta_{t}$ (8)
+Under the Assumption 11, $\epsilon_{\theta}$ is able to accuaratly predict the noise added on the $\sqrt{\bar{\alpha}_{t}x_{0}}$ to obtain $x_{t}$, hence the prediction of $\epsilon_{\theta}$ in Equation 7 should be $\rho\epsilon_{t}$.
+We have by substituting $\rho\epsilon_{t}$ into Equation 7:
+$\mu_{\theta}(x^{\prime}_{t},t)=\frac{1}{\sqrt{\alpha_{t}}}\Big{(}x^{\prime}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho}\epsilon_{t}\Big{)}$ (9)
+By comparing Equation 9 and Equation 8 with Equation 8 , we found the mean of the reverse process is the same when inputting the $x^{\prime}_{T}$ to the $f_{\theta}$ and $f_{\tilde{\theta}}$, while the variance of $f_{\tilde{\theta}}$ is $\rho^{2}$ times larger than $f_{\theta}$.
+For simplicity, Let $a_{t}=\mu_{\theta}(x^{\prime}_{t},t)$ and $b_{t}=\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\cdot\beta_{t}$, we have: $q_{\theta}(x_{t-1}|x_{t})\sim\mathcal{N}(a_{t},b_{t}I)$ and $q_{\tilde{{\theta}}}(x_{t-1}|x_{t})\sim\mathcal{N}(a_{t},b_{t}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho^{2}}I)$.
+Hence, by the reparameterization trick, the variance of the generated $\hat{\tilde{x_{0}}}$ of $f_{\tilde{\theta}}$ is $\rho^{2}$ times greater than $f_{\theta}$.
+Without loss of generality, we use the Gaussian distribution to discribe the output distribution.
+Given the Assumption 11, and $q(x)\sim\mathcal{N}(x_{c},\sigma_{c}I)$, $\hat{x_{0}}=f_{\theta}(x^{\prime}_{T})\sim\mathcal{N}(x_{c},\frac{\sigma_{c}}{\rho^{2}}I)$, which completes the proof.
+∎
+Appendix P Proof of Theorem 2
+Theorem 8.
+Suppose the output domain of the diffusion model $f_{\theta}$ is Gaussian.
+Let $\mathcal{N}(\mu_{c},\sigma_{c})$ and $\mathcal{N}(\mu_{b},\sigma_{b})$ denote the distribution of clean generations and backdoor generations respectively.
+Let $N$ denote the image size.
+We assume that $\sigma_{c}\geq\sigma_{b}+\rho^{2}$.
+Given clean input noise $x_{T}^{c}\sim\mathcal{N}(0,I)$, backdoor input noise $x_{T}^{b}=x_{T}^{c}+\delta\sim\mathcal{N}(\delta,I)$, and Lemma 1 if we perturb the clean input noises $x_{T}$ and backdoor input noise $x_{T}^{b}$ with some $\epsilon\sim\mathcal{N}(0,I)$ simultaneously, then for the resulted clean generations $\mathcal{N}(\mu^{\prime}_{c},\sigma^{\prime}_{c})$ and the backdoor generations $\mathcal{N}(\mu^{\prime}_{b},\sigma^{\prime}_{b})$, we have that $\sigma^{\prime}_{c}-\sigma^{\prime}_{b}\geq 1$.
 Proof.
 We begin our proof by first introducing the basic diffusion process for clean samples.
-Definition G.2 (Clean Forward process).
+Definition 9 (Clean Forward process).
 Let $x_{0}\sim q(x)$ denote a sample from the clean data distribution, $x_{T}\sim\mathcal{N}(0,I)$ denote the pure Gaussian noise.
-Given the variance schedule $\{\beta_{t}\}_{t=1}^{T}$ in DDPM [22], define the forward process to diffuse $x_{0}$ to $x_{T}$ for clean samples:
+Given the variance schedule $\{\beta_{t}\}_{t=1}^{T}$ in DDPM (Ho, Jain, and Abbeel 2020), define the forward process to diffuse $x_{0}$ to $x_{T}$ for clean samples:
 $q(x_{t}|x_{t-1})=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}I)$ (10)
@@
 Here, we give the definition of the reverse process of clean samples:
-Definition G.3 (Clean Reverse process).
+Definition 10 (Clean Reverse process).
 The reverse process for clean samples is
@@
 $\Sigma_{\theta}(x_{t},t)=\frac{(1-\bar{\alpha}_{t-1})\beta_{t}}{1-\bar{\alpha}_{t}},$ (14)
-Detailed proof can be found in [24]. In our setting, we assume that the attacker is able to deploy a well-trained diffusion model on the internet. Accordingly, we make the following assumptions:
-Assumption G.4.
+Detailed proof can be found in (Hu et al.
+2023).
+In our setting, we assume that the attacker is able to deploy a well-trained diffusion model on the internet.
+Accordingly, we make the following assumptions:
+Assumption 11.
 Assume a well-trained clean diffusion model $f_{\theta}$, designed to generate clean samples $x_{o}\sim q(x)$ from pure Gaussian noise $x_{T}\sim\mathcal{N}(0,I)$.
@@
 As a result, both $f_{\theta}$ and $f_{\tilde{\theta}}$ can generate images following clean data distribution $q(x)$, given inputs following $\mathcal{N}(0,I)$ and $\mathcal{N}(0,\rho^{2}I)$, respectively.
-The forward and backward processes of $f_{\theta}$ are already defined from Equation 3 to 7.
+The forward and backward processes of $f_{\theta}$ are already defined from Equation 10 to 14.
 Notably, in our analysis, $\rho^{2}$ is set to 2 for clean samples to account for the addition of Gaussian noise.
@@
 Then the $f_{\tilde{\theta}}$ aims to learn a generative process, such that $p_{\tilde{\theta}}(x^{\prime}_{t-1}|x^{\prime}_{t})=q(x^{\prime}_{t-1}|x^{\prime}_{t})$, which is,
-$\begin{split}q(x^{\prime}_{t-1}|x^{\prime}_{t},x_{0})&=q(x^{\prime}_{t}|x^{\prime}_{t-1},x_{0})\frac{q(x^{\prime}_{t-1}|x_{0})}{q(x^{\prime}_{t}|x_{0})}\\ &\propto\exp\Big{(}-\frac{1}{2}\big{(}\frac{(x^{\prime}_{t}-\sqrt{\alpha_{t}}x^{\prime}_{t-1})^{2}}{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho^{2}}\beta_{t}}+\frac{(x^{\prime}_{t-1}-\sqrt{\bar{\alpha}_{t-1}}x_{0})^{2}}{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho^{2}}(1-\bar{\alpha}_{t-1})}-\frac{(x^{\prime}_{t}-\sqrt{\bar{\alpha}_{t}}x_{0})^{2}}{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho^{2}}(1-\bar{\alpha}_{t})}\big{)}\Big{)}\\ &=\exp\Big{(}-\frac{1}{2}\big{(}\frac{{x^{\prime}}_{t}^{2}-2\sqrt{\alpha_{t}}x_{t}x_{t-1}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{+\alpha_{t}}{x^{\prime}}_{t-1}^{2}}{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho^{2}}\beta_{t}}+\frac{{x^{\prime}}_{t-1}^{2}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{-2\sqrt{\bar{\alpha}_{t-1}}x_{0}}x^{\prime}_{t-1}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{+\bar{\alpha}_{t-1}x_{0}^{2}}}{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho^{2}}(1-\bar{\alpha}_{t-1})}-\frac{(x^{\prime}_{t}-\sqrt{\bar{\alpha}_{t}}x_{0})^{2}}{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho^{2}}(1-\bar{\alpha}_{t})}\big{)}\Big{)}\\ &=\exp\Big{(}-\frac{1}{2{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho^{2}}}\big{(}(\frac{\alpha_{t}}{\beta_{t}}+\frac{1}{1-\bar{\alpha}_{t-1}}){x^{\prime}}_{t-1}^{2}-(\frac{2\sqrt{\alpha_{t}}}{\beta_{t}}x^{\prime}_{t}+\frac{2\sqrt{\bar{\alpha}_{t-1}}}{1-\bar{\alpha}_{t-1}}x_{0})x^{\prime}_{t-1}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{+C(x^{\prime}_{t},x^{\prime}_{0})\big{)}\Big{)}}\\ &:=\mathcal{N}(x^{\prime}_{t-1};\mu_{\tilde{\theta}}(x^{\prime}_{t},t),\Sigma_{\tilde{\theta}}(x^{\prime}_{t},t)),\end{split}$ (10)
+$\begin{split}q(x^{\prime}_{t-1}|x^{\prime}_{t},x_{0})&=q(x^{\prime}_{t}|x^{\prime}_{t-1},x_{0})\frac{q(x^{\prime}_{t-1}|x_{0})}{q(x^{\prime}_{t}|x_{0})}\\ &\propto\exp\Big{(}-\frac{1}{2}\big{(}\frac{(x^{\prime}_{t}-\sqrt{\alpha_{t}}x^{\prime}_{t-1})^{2}}{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho^{2}}\beta_{t}}+\frac{(x^{\prime}_{t-1}-\sqrt{\bar{\alpha}_{t-1}}x_{0})^{2}}{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho^{2}}(1-\bar{\alpha}_{t-1})}-\frac{(x^{\prime}_{t}-\sqrt{\bar{\alpha}_{t}}x_{0})^{2}}{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho^{2}}(1-\bar{\alpha}_{t})}\big{)}\Big{)}\\ &=\exp\Big{(}-\frac{1}{2}\big{(}\frac{{x^{\prime}}_{t}^{2}-2\sqrt{\alpha_{t}}x_{t}x_{t-1}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{+\alpha_{t}}{x^{\prime}}_{t-1}^{2}}{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho^{2}}\beta_{t}}+\frac{{x^{\prime}}_{t-1}^{2}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{-2\sqrt{\bar{\alpha}_{t-1}}x_{0}}x^{\prime}_{t-1}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{+\bar{\alpha}_{t-1}x_{0}^{2}}}{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho^{2}}(1-\bar{\alpha}_{t-1})}-\frac{(x^{\prime}_{t}-\sqrt{\bar{\alpha}_{t}}x_{0})^{2}}{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho^{2}}(1-\bar{\alpha}_{t})}\big{)}\Big{)}\\ &=\exp\Big{(}-\frac{1}{2{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho^{2}}}\big{(}(\frac{\alpha_{t}}{\beta_{t}}+\frac{1}{1-\bar{\alpha}_{t-1}}){x^{\prime}}_{t-1}^{2}-(\frac{2\sqrt{\alpha_{t}}}{\beta_{t}}x^{\prime}_{t}+\frac{2\sqrt{\bar{\alpha}_{t-1}}}{1-\bar{\alpha}_{t-1}}x_{0})x^{\prime}_{t-1}\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{+C(x^{\prime}_{t},x^{\prime}_{0})\big{)}\Big{)}}\\ &:=\mathcal{N}(x^{\prime}_{t-1};\mu_{\tilde{\theta}}(x^{\prime}_{t},t),\Sigma_{\tilde{\theta}}(x^{\prime}_{t},t)),\end{split}$
 Following the standard Gaussian density function, the mean and variance can be parameterized as follows.
-$\begin{split}\Sigma_{\tilde{\theta}}(x^{\prime}_{t},t)&=1/{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho^{2}}(\frac{\alpha_{t}}{\beta_{t}}+\frac{1}{1-\bar{\alpha}_{t-1}})=1/(\frac{\alpha_{t}-\bar{\alpha}_{t}+\beta_{t}}{\beta_{t}(1-\bar{\alpha}_{t-1})})=\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\cdot\beta_{t}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho^{2}}\\ \mu_{\tilde{\theta}}(x^{\prime}_{t},t)&=\frac{1}{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho^{2}}}(\frac{\sqrt{\alpha_{t}}}{\beta_{t}}x^{\prime}_{t}+\frac{\sqrt{\bar{\alpha}_{t-1}}}{1-\bar{\alpha}_{t-1}}x_{0})/\frac{1}{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho^{2}}}(\frac{\alpha_{t}}{\beta_{t}}+\frac{1}{1-\bar{\alpha}_{t-1}})\\ &=(\frac{\sqrt{\alpha_{t}}}{\beta_{t}}x^{\prime}_{t}+\frac{\sqrt{\bar{\alpha}_{t-1}}}{1-\bar{\alpha}_{t-1}}x_{0})\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\cdot\beta_{t}\\ &=\frac{\sqrt{\alpha_{t}}(1-\bar{\alpha}_{t-1})}{1-\bar{\alpha}_{t}}x^{\prime}_{t}+\frac{\sqrt{\bar{\alpha}_{t-1}}\beta_{t}}{1-\bar{\alpha}_{t}}x_{0}\\ &=\frac{\sqrt{\alpha_{t}}(1-\bar{\alpha}_{t-1})}{1-\bar{\alpha}_{t}}x^{\prime}_{t}+\frac{\sqrt{\bar{\alpha}_{t-1}}\beta_{t}}{1-\bar{\alpha}_{t}}\frac{1}{\sqrt{\bar{\alpha}_{t}}}(x^{\prime}_{t}-\sqrt{1-\bar{\alpha}_{t}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho}\boldsymbol{\epsilon}_{t})\\ &=\frac{1}{\sqrt{\alpha_{t}}}\Big{(}x^{\prime}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho}\boldsymbol{\epsilon}_{t}\Big{)}\end{split}$ (11)
-According to Lemma 3.2, if we add Gaussian noise to the origin input image, which results in $\mathcal{N}(0,\rho^{2}I)$, then the distribution of generated images of the diffusion model has the same mean, but a variance scaled by $\frac{1}{\rho^{2}}$, where $\rho^{2}=2$ for clean samples.
+$\begin{split}\Sigma_{\tilde{\theta}}(x^{\prime}_{t},t)&=1/{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho^{2}}(\frac{\alpha_{t}}{\beta_{t}}+\frac{1}{1-\bar{\alpha}_{t-1}})=1/(\frac{\alpha_{t}-\bar{\alpha}_{t}+\beta_{t}}{\beta_{t}(1-\bar{\alpha}_{t-1})})=\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\cdot\beta_{t}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho^{2}}\\ \mu_{\tilde{\theta}}(x^{\prime}_{t},t)&=\frac{1}{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho^{2}}}(\frac{\sqrt{\alpha_{t}}}{\beta_{t}}x^{\prime}_{t}+\frac{\sqrt{\bar{\alpha}_{t-1}}}{1-\bar{\alpha}_{t-1}}x_{0})/\frac{1}{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho^{2}}}(\frac{\alpha_{t}}{\beta_{t}}+\frac{1}{1-\bar{\alpha}_{t-1}})\\ &=(\frac{\sqrt{\alpha_{t}}}{\beta_{t}}x^{\prime}_{t}+\frac{\sqrt{\bar{\alpha}_{t-1}}}{1-\bar{\alpha}_{t-1}}x_{0})\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\cdot\beta_{t}\\ &=\frac{\sqrt{\alpha_{t}}(1-\bar{\alpha}_{t-1})}{1-\bar{\alpha}_{t}}x^{\prime}_{t}+\frac{\sqrt{\bar{\alpha}_{t-1}}\beta_{t}}{1-\bar{\alpha}_{t}}x_{0}\\ &=\frac{\sqrt{\alpha_{t}}(1-\bar{\alpha}_{t-1})}{1-\bar{\alpha}_{t}}x^{\prime}_{t}+\frac{\sqrt{\bar{\alpha}_{t-1}}\beta_{t}}{1-\bar{\alpha}_{t}}\frac{1}{\sqrt{\bar{\alpha}_{t}}}(x^{\prime}_{t}-\sqrt{1-\bar{\alpha}_{t}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho}\boldsymbol{\epsilon}_{t})\\ &=\frac{1}{\sqrt{\alpha_{t}}}\Big{(}x^{\prime}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho}\boldsymbol{\epsilon}_{t}\Big{)}\end{split}$
+According to Lemma 1, if we add Gaussian noise to the origin input image, which results in $\mathcal{N}(0,\rho^{2}I)$, then the distribution of generated images of the diffusion model has the same mean, but a variance scaled by $\frac{1}{\rho^{2}}$, where $\rho^{2}=2$ for clean samples.
 Now we start analyzing the backdoor samples.
-Definition G.5 (Backdoor Forward process).
+Definition 12 (Backdoor Forward process).
 Let $x_{0}^{b}\sim q(x_{b})$ denote a sample from target data distribution, $\delta$ denote a trigger, and $x_{T}^{b}\sim\mathcal{N}(\delta,I)$ denote the pure Gaussian noise attached by a trigger.
-Given the variance schedule $\{\beta_{t}\}_{t=1}^{T}$ in DDPM [22], define the forward process to diffuse $x_{0}^{b}$ to $x_{T}^{b}$ for backdoor samples:
+Given the variance schedule $\{\beta_{t}\}_{t=1}^{T}$ in DDPM (Ho, Jain, and Abbeel 2020), define the forward process to diffuse $x_{0}^{b}$ to $x_{T}^{b}$ for backdoor samples:
 $q(x_{t}^{b}|x_{t-1}^{b})=\mathcal{N}(x_{t}^{b};\sqrt{1-\beta_{t}}x_{t-1_{b}}+k_{t}\delta,\beta_{t}I),$ (17)
@@
 Specifically, we consider that the clean data distribution $q(x)$ of the clean samples follow $\mathcal{N}(x_{c},\sigma_{c}I)$, while the backdoor samples follow $\mathcal{N}(x_{b},\sigma_{b}I)$.
-For D2I attack, the target generated image is specific (e.g., Mickey Mouse), thus its variance is 0.
-For In-D2D and Out-D2D, the target images belong to the same class, hence their variance should be much lower than that of images generated from the clean model, as those are from various, completely different classes.
-Therefore, under the above case and Lemma 3.2, the distributions generated by clean and backdoor samples after noise addition are $\mathcal{N}(x_{c},\sigma_{c}\frac{1}{\rho^{2}}I)$ and $\mathcal{N}(x_{b},\sigma_{b}\frac{1}{\rho^{2}}I)$, respectively.
-In particular, we consider three attack methods.
-Hence, we have three kinds of $\sigma_{b}$.
-Let $\sigma_{D2I}$ denote the variance of D2I attack, $\sigma_{iD2D}$ denote the variance of In-D2I attack, $\sigma_{oD2I}$ denote the variance of Out-D2I attack defined in Appendix  B.
-We then have,
-$0=\frac{1}{\rho^{2}}\sigma_{D2I}\ll\frac{1}{\rho^{2}}\sigma_{In-D2D},\frac{1}{\rho^{2}}\sigma_{Out-D2D}\ll\frac{1}{\rho^{2}}\sigma_{c}.$ (14)
-Equation 14 shows the different variances of the images generated from the fixed diffusion model $f_{\theta}$ with noise added to the input in terms of clean and backdoor samples.
-Under Lemma 3.1, which states that two points sample from a distribution of larger variance tend to have a larger distance than those from a distribution of lower variance, we have that after noise addition the backdoor inputs tend to generate similar images, while the clean inputs tend to generate diversified images.
-Take TrojDiff(In-D2D) attack on Cifar10 for example, $N$ is chosen as $32\times 32\times 3=3072$.
-Based on Lemma 3.1, we must have that $\|x_{1}-x_{2}\|_{2}\geq\|x_{3}-x_{4}\|_{2}$ under the condition of $\sigma_{In-D2D}<\frac{3071}{3072}\sigma_{c}$, where $x_{1},x_{2}\sim\mathcal{N}(x_{c},\sigma_{c}\frac{1}{\rho^{2}}I)$ and $x_{3},x_{4}\sim\mathcal{N}(x_{b},\sigma_{In-D2D}\frac{1}{\rho^{2}}I)$.
-However, according to Equation 14, this conditional is trivial to achieve, since $\frac{1}{\rho^{2}}\sigma_{c}\gg\frac{1}{\rho^{2}}\sigma_{In-D2D}$.
+Therefore, under the Lemma 1, the distributions generated by clean and backdoor samples after noise addition are $\mathcal{N}(x_{c},\sigma^{\prime}_{c}I)$ and $\mathcal{N}(x_{b},\sigma_{b}^{\prime}I)$, respectively, where $\sigma^{\prime}_{c}=\sigma_{c}\frac{1}{\rho^{2}}$ and $\sigma_{b}^{\prime}=\sigma_{b}\frac{1}{\rho^{2}}$.
+In reality, the variance of clean generations $\sigma_{c}$ can be a much larger value than that of the backdoor generations.
+Based on the assumption that $\sigma_{c}-\sigma_{b}>\rho^{2}$, we then have that $\sigma_{c}^{\prime}>\sigma_{b}^{\prime}+1$.
 This completes the proof.
-Appendix H Proof of Lemma 3.2
-Lemma H.1.
-Let $f_{\theta}$ and $f_{\tilde{\theta}}$ be two well-trained diffusion models as defined in the Assumption G.4.
-Let $x^{\prime}_{T}$ follow $\mathcal{N}(0,\rho^{2}I)$.
-Let the clean data distribution be $q(x)\sim\mathcal{N}(x_{c},\sigma_{c}I)$.
-We then have:
-$\hat{x_{0}}=f_{\theta}(x^{\prime}_{T})\sim\mathcal{N}(x_{c},\frac{\sigma_{c}}{\rho^{2}}I)$ (15)
+Lemma 13 (Bounds on Expected Length of Gaussian Random Variable (Chandrasekaran et al. 2012)).
+Given that $x\sim\mathcal{N}(\mu,\sigma I)$, where $x$ is a $N$-dimensional vector. Let $\mathbb{E}(X)$ be the expectation value of the random variable $X$. Then, we have
+$\frac{N}{\sqrt{N+1}}\leq\sigma^{-1}\mathbb{E}(\|x\|_{2})\leq\sqrt{N}$
+Corollary 14.
+[…]
+$\mathbb{E}(\|x_{1}-x_{2}\|_{2}-\|x_{3}-x_{4}\|)\geq\frac{N(\sigma_{c}-\sigma_{b})-\sigma_{b}}{\sqrt{N+1}}>0.$
 Proof.
-The output generated image from $f_{\tilde{\theta}}$ when input $x^{\prime}_{T}$ is given follows:
-$\hat{\tilde{x_{0}}}=f_{\tilde{\theta}}(x^{\prime}_{T})\sim\mathcal{N}(x_{c},\sigma_{c}I),$ (16)
-the Equation 16 is due to Assumption G.4.
-In particular, to obtain the generated image follows $q(x)$, the reverse process is defined as $q(x^{\prime}_{t-1}|x^{\prime}_{t})\sim\mathcal{N}(x^{\prime}_{t-1};\mu_{\tilde{\theta}}(x^{\prime}_{t},t),\Sigma_{\tilde{\theta}}(x^{\prime}_{t},t))$, where $\mu_{\tilde{\theta}}(x^{\prime}_{t},t)=\frac{1}{\sqrt{\alpha_{t}}}\Big{(}x^{\prime}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho}{\epsilon}_{t}\Big{)}$ and $\Sigma_{\tilde{\theta}}(x^{\prime}_{t},t)=\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\cdot\beta_{t}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho^{2}}$ (Equation 11).
-For the $f_{\theta}$, although the reverse process is also a gaussian distribution, $\mu_{\theta}(x_{t},t)=\frac{1}{\sqrt{\alpha_{t}}}\Big{(}x_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\boldsymbol{\epsilon}_{\theta}(x_{t},t)\Big{)}$, $\Sigma_{\theta}(x_{t},t)=\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\cdot\beta_{t}$ (Equation 6 and 7).
-To obtain the generated image from $f_{\theta}$ when input $x^{\prime}_{T}$ is given, we substitute $x_{t}=x^{\prime}_{t}$ into the fixed $f_{\theta}$. We then have by Equation 9 that:
-$\mu_{\theta}(x^{\prime}_{t},t)=\frac{1}{\sqrt{\alpha_{t}}}\Big{(}x^{\prime}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\boldsymbol{\epsilon}_{\theta}(x^{\prime}_{t}=\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho}\epsilon,t)\Big{)}$ (17)
-$\Sigma_{\theta}(x^{\prime}_{t},t)=\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\cdot\beta_{t}$ (18)
-Under the Assumption G.4, $\epsilon_{\theta}$ is able to accuaratly predict the noise added on the $\sqrt{\bar{\alpha}_{t}x_{0}}$ to obtain $x_{t}$, hence the prediction of $\epsilon_{\theta}$ in Equation 17 should be $\rho\epsilon_{t}$.
-We have by substituting $\rho\epsilon_{t}$ into Equation 17:
-$\mu_{\theta}(x^{\prime}_{t},t)=\frac{1}{\sqrt{\alpha_{t}}}\Big{(}x^{\prime}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho}\epsilon_{t}\Big{)}$ (19)
-By comparing Equation 19 and Equation 18 with Equation 11 , we found the mean of the reverse process is the same when inputting the $x^{\prime}_{T}$ to the $f_{\theta}$ and $f_{\tilde{\theta}}$, while the variance of $f_{\tilde{\theta}}$ is $\rho^{2}$ times larger than $f_{\theta}$.
-For simplicity, Let $a_{t}=\mu_{\theta}(x^{\prime}_{t},t)$ and $b_{t}=\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}\cdot\beta_{t}$, we have: $q_{\theta}(x_{t-1}|x_{t})\sim\mathcal{N}(a_{t},b_{t}I)$ and $q_{\tilde{{\theta}}}(x_{t-1}|x_{t})\sim\mathcal{N}(a_{t},b_{t}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rho^{2}}I)$.
-Hence, by the reparameterization trick, the variance of the generated $\hat{\tilde{x_{0}}}$ of $f_{\tilde{\theta}}$ is $\rho^{2}$ times greater than $f_{\theta}$.
-Without loss of generality, we use the Gaussian distribution to discribe the output distribution.
-Given the Assumption G.4, and $q(x)\sim\mathcal{N}(x_{c},\sigma_{c}I)$, $\hat{x_{0}}=f_{\theta}(x^{\prime}_{T})\sim\mathcal{N}(x_{c},\frac{\sigma_{c}}{\rho^{2}}I)$, which completes the proof.
-∎
-Appendix I Proof of Lemma 3.1
-Lemma I.1 [10]).
-Given that $x\sim\mathcal{N}(\mu,\sigma I)$, where $x$ is a $N$-dimensional vector.
-Then, we must have $\frac{N}{\sqrt{N+1}}\leq\sigma^{-1}\mathbb{E}(\|x\|_{2})\leq\sqrt{N}$, where $\mathbb{E}(X)$ denotes an expectation value of the random variable $X$.
-Lemma I.2.
-[…]
-Proof.
-Let $A:=\|x_{1}-x_{2}\|_{2}$, $B:=\|x_{3}-x_{4}\|_{2}$, $C:=B-A$, and $t$ be some arbitrary positive number. According to the Markov inequality, we have that
-$\begin{split}Pr(C>t)\leq\frac{\mathbb{E}(C)}{t}=\frac{\mathbb{E}(B-A)}{t}=\frac{\mathbb{E}(B)-\mathbb{E}(A)}{t}=\frac{\mathbb{E}(\|x_{3}-x_{4}\|_{2})-\mathbb{E}(\|x_{1}-x_{2}\|_{2})}{t}\end{split}$ (20)
-By basic calculations, it is obvious that $x_{1}-x_{2}\sim\mathcal{N}(0,\sqrt{2}\sigma_{1})$ and $x_{3}-x_{4}\sim\mathcal{N}(0,\sqrt{2}\sigma_{2})$. Then, according to Lemma I.1, we obtain that,
-$Pr(C>t)\leq\sqrt{2}(\frac{\sigma_{2}\sqrt{N}}{t}-\frac{N\sigma_{1}}{t\sqrt{N+1}})\leq\sqrt{2}\frac{\sigma_{2}(N+1)-N\sigma_{1}}{t\sqrt{N+1}}$ (21)
-If we have that $\sigma_{2}<\frac{N}{N+1}\sigma_{1}$, then the numerator will be always less than 0. This completes the proof. ∎
+It is obvious that $x_{1}-x_{2}\sim\mathcal{N}(0,\sqrt{2}\sigma_{c})$ and $x_{3}-x_{4}\sim\mathcal{N}(0,\sqrt{2}\sigma_{b})$. Then, According to Lemma 13, we obtain that
+$\begin{split}\mathbb{E}(\|x_{1}-x_{2}\|_{2}-\|x_{3}-x_{4}\|)&=\mathbb{E}(\|x_{1}-x_{2}\|_{2})-\mathbb{E}(\|x_{3}-x_{4}\|)\\ &\geq\sqrt{2}(\frac{N\sigma_{c}}{\sqrt{N+1}}-\sqrt{N}\sigma_{b})\\ &=\sqrt{2}\frac{N\sigma_{c}-\sqrt{N(N+1)}\sigma_{b}}{\sqrt{N+1}}\\ &\geq\sqrt{2}\frac{N\sigma_{c}-(N+1)\sigma_{b}}{\sqrt{N+1}}\\ &=\sqrt{2}\frac{N(\sigma_{c}-\sigma_{b})-\sigma_{b}}{\sqrt{N+1}}\end{split}$ (19)
+Based on the conclusion from the Theorem 2, we have that $\sigma_{c}-\sigma_{b}>1$. Therefore, we obtain that,
+$\mathbb{E}(\|x_{1}-x_{2}\|_{2}-\|x_{3}-x_{4}\|)>\sqrt{2}\frac{N-\sigma_{b}}{\sqrt{N+1}}\geq 0,$ (20)
+where the last inequality hold because $\sigma_{b}=\mathcal{O}(N)$. This completes the proof. ∎